### v1

# ChartBlender: An Interactive System for Authoring and Synchronizing Visualization Charts in Video

Abstract
Embedding data visualizations in video can enhance the communication of complex information. However, this process is often labor-intensive, requiring designers to adjust visualizations frame by frame manually. In this work, we present ChartBlender, a novel system that streamlines this process by enabling users to create data visualizations, embed them seamlessly into video scenes, and automatically synchronize them with both camera motion and moving objects. Particularly, ChartBlender incorporates a tracking algorithm that supports both object and camera tracking, ensuring robust alignment of visualizations with dynamic video content. To maintain visual clarity and aesthetic coherence, we also explore the design space of video-suited visualizations and develop a library of customizable templates optimized for video embedding. We evaluate ChartBlender through two controlled experiments and expert interviews with five domain experts. Results show that our system enables accurate synchronization and accelerates the production of data-driven videos.
Index Terms:
data visualization, video editing, object tracking, camera tracking, interface, visual storytelling.
publicationid: pubid: $0$–$00$ 2021 IEEE
1 Introduction
Embedding data visualizations into videos has emerged as a powerful method for conveying complex information [1, 2, 3, 4, 5]. Such a form of representation is increasingly used across various domains (e.g., sports, traffic, and journalism), allowing audiences to access supplementary information without being distracted from the primary video content[6]. Despite its potential, the process remains labor-intensive: achieving spatial and temporal synchronization between visualizations and video typically requires manual, frame-by-frame alignment. Even a 10-second visualization-embedded video may demand up to a full day of manual work from designers.
Motivated by a collaborative project with a major Chinese media organization to create a chart-embedded video, we conducted a comprehensive survey of existing tools and systems. On one end of the spectrum are professional tools like Adobe After Effects [7] and Blender[8]. While powerful, these tools rely on manual parameter tuning and lack intelligent assistance, making them difficult to use efficiently. On the other end, recent advances in Augmented Reality (AR) and Virtual Reality (VR) have inspired research systems that allow casual users to overlay visualizations on videos for sharing quickly [9]. These systems prioritize ease of use and rapid output but are designed primarily for end users with simple needs. As a result, they fall short in supporting high-quality, professionally produced content. Specifically, they lack robust motion tracking for objects with diverse movement patterns, as well as precise control over visualization placement and editing capabilities essential for content creators and news organizations.
Therefore, there is a pressing need for an interactive authoring system supports flexible, efficient embedding of data visualizations into complex video scenes and ensures accurate motion synchronization throughout the video. However, to design such an effective system, three challenges need to be tackled: (1) track the motion of videos accounting for both moving objects in the scene and camera pose variations, to ensure consistent synchronization between visualizations and the video scene; (2) design generalized visualizations to support clear and informative presentation of data within dynamic video contexts; and (3) enable user-driven authoring and refinement of visualizations through simple and integrated interaction. Addressing these challenges is essential for the precise and seamless embedding of data visualizations into real-world videos.
To address these challenges, we introduce ChartBlender, a novel authoring system for embedding and automatically synchronizing data visualizations with video during playback. In particular, the user simply selects a target tracking point and a preferred tracking mode, and the system automatically generates the complete motion trajectory for visualization throughout the Users can edit the visualization at any time, including both its visual style and its 3D pose within the video scene. In this way, the system supports the human-machine collaboration to generate data-driven videos rapidly while ensuring the visualization appears visually coherent and seamlessly embedded into the video. The major contributions of the paper are as follows:
  • System. We introduce ChartBlender, a novel authoring system that enables the authoring and automatic synchronization of data visualizations with video content. It features an interactive interface that supports the end-to-end creation of visualization-embedded videos through an integrated authoring workflow.
  • Design Space. We characterize the design space of visualizations suitable for video embedding. Specifically, we identify two primary motion modes, camera tracking and object tracking, and identify five commonly used chart types. Based on this analysis, we develop a library of generalized visualization templates designed to convey complex data clearly and effectively within dynamic video contexts.
  • Algorithms. We develop two algorithms to support robust visualization embedding in dynamic video scenes. The first estimates frame-by-frame camera poses, enabling visualizations to remain stably anchored to the background. The second computes and refines object trajectories across frames, allowing visualizations to accurately follow moving targets within the video.
2 Related Work
In this section, we survey the papers that are most relevant to our work, including professional video editing software, embedded visualization, and point tracking.
2.1 Video Editing Software
Video editing is a broad field that involves various techniques, technologies, and tools used to create, manipulate, and refine video content, which often pose barriers to users [10, 11]. Professional video editing software offers powerful tools to insert animated objects into videos. Traditionally, users must manually keyframe an object’s position, rotation, and scale over time, with smooth transitions achieved through interpolation. While recent advances have introduced automated tracking methods, such as 3D camera estimation and object tracking, these techniques, available in tools like Adobe After Effects, Blender, and Nuke [7, 8, 12], still require a level of expertise that can be challenging for beginners. At the same time, the growing demand for data-driven videos highlights a gap in current workflows: mainstream video editing tools lack integrated support for data visualization. Users often switch between separate tools or project files to create visualizations, making it difficult to assess whether a chart fits naturally within a given video context. Currently, no tool allows for direct editing of chart styles within the video environment while supporting real-time assessment, motion synchronization, and object tracking. This creates a pressing need for authoring tools that bridge the gap between visualization and video editing.
2.2 Embedded Visualization in Video
Embedded data visualization in videos has become increasingly popular [13, 14, 15, 16]. However, this process often requires users to manually create keyframes for positioning the visualizations within the video. Recently, there has been increasing research on intelligent systems that can help embed visualizations in videos [17, 18, 6, 19]. Early works primarily focused on embedding static charts into videos. For instance, SmartShot [17] developed a system that optimizes the placement of charts to avoid obstructing the visual focus, ensuring aesthetic balance and readability. SwimFlow [18] explored how to embed visualizations into swimming competition videos. Their system allows charts to animate by using manual annotations to mark athletes’ starting positions and interpolating to calculate their motion trajectories. However, this approach is clearly limited when dealing with more complex motion patterns. Recent advancements in computer vision now enable the embedding of visualizations directly into sports videos, offering a more dynamic and adaptive solution. For instance, iBall [6] embeds dynamic visual data cues into basketball videos and automatically adjusts the information display based on the user’s gaze. However, when the camera moves, the charts can only approximate the motion rather than being stably anchored to the scene. Additionally, these methods typically involve 2D overlays and tracking.
In contrast to these domain-specific solutions, our approach tracks 3D trajectories of arbitrary objects and supports 3D effects, including perspective transformations and occlusions of the inserted visualizations. Moreover, existing systems often lack the flexibility to freely create and edit visualizations to meet user-specific needs. In our system, users can directly edit the types, styles, and data mappings of charts on the video, streamlining the creation process and offering greater customization.
2.3 Point Tracking
Tracking arbitrary points in videos has witnessed substantial progress in recent years [20, 21, 22], with methods such as TAP-Net [23] demonstrating the ability to reliably follow both static and dynamic points throughout a video sequence. These approaches typically leverage appearance-based cues and temporal consistency to maintain accurate correspondences, even under significant motion and occlusion. However, focusing solely on 2D tracking is often insufficient in practical applications. While a dynamic point in the scene can be tracked frame-by-frame, certain use cases, such as visual overlays or interactive annotations, require this point to remain anchored relative to the static scene background, rather than to moving objects or transient visual cues. To achieve this, it becomes necessary to disentangle camera motion from scene dynamics, which calls for robust camera pose estimation.
Fortunately, a variety of mature techniques now exist for estimating camera pose from monocular or multi-view video streams [24, 25], ranging from classical SLAM-based pipelines [26, 27, 28] to deep learning methods[29, 30, 31, 32, 33]. The integration of 2D point tracking with camera pose estimation enables a more unified paradigm: tracking arbitrary points in 3D space. This formulation has recently attracted increasing attention [34, 35, 36], as researchers extend traditional 2D tracking frameworks into the 3D domain. For instance, SpatialTracker[36] is the first work specifically designed for the task of 3D point tracking. It estimates the 3D trajectory of arbitrary pixels by combining image features with monocular depth estimation, and propagates information through a graph neural network (GNN) to model long-range spatiotemporal motion. Similarly, TAPIR-3D[34] is an experimental extension of the original 2D point tracking model TAPIR, designed to output 3D point trajectories. It initializes keypoints in each frame, refines their 2D tracks using temporal context from adjacent frames, and subsequently lifts them to 3D space using depth estimation models such as ZoeDepth [37].
We propose a novel approach that enhances the accuracy of 3D point tracking by integrating it with the practical requirements of embedding visualizations into real-world video content. Our system enables users to anchor visualizations to fixed locations within the scene or dynamically attach them to moving objects, regardless of camera motion or scene complexity. This flexibility allows for robust visualization embedding across a wide range of video scenarios.
3 System Design and Overview
The ChartBlender system was motivated by our work experience on a data video design in collaboration with a prominent Chinese media organization. In this project, we were responsible for designing visualization charts and embedding the charts into a video. To support this task, we surveyed existing tools and systems for creating chart-embedded video. One notable system was VisTellAR [9], which enables end-users to record short videos and embed data visualizations using augmented reality (AR) techniques. VisTellAR tracks static elements in the video and anchors charts accordingly, offering a lightweight and mobile-friendly approach.
Although VisTellAR is not open-source, we introduced it to our collaborators as a representative example of chart-embedded video tools. They noted, however, that end-users seldom embed charts into videos directly. In contrast, professional data journalists (our target users) require such capabilities but demand far greater precision and control. They highlighted three critical requirements: (1) accurate video motion tracking that can handle both static and dynamic objects; (2) compatibility with diverse video sources, including those lacking camera metadata, as is common in real-world media production; and (3) support for multiple dynamic charts within a single video, with fine-grained control over timing and placement to maintain contextual relevance and visual coherence.
We observed a significant gap in tools that enable seamless and efficient embedding of visualizations into real-world videos. In current practice, designers often resort to manually inserting pre-rendered charts frame by frame using software like Adobe After Effects. This process so labor-intensive that a ten-second clip can require an entire day of work according to our own experience (see the example video  [^1] http://edu.people.com.cn/n1/2023/0915/c1006-40078200.html (2’44”, 3’25”) that we produced in this way). The majority of this effort goes into repetitive adjustments of position, scale, and orientation, rather than creative or analytical tasks. This inefficiency underscores the need for a dedicated authoring tool that can automate chart-video synchronization. Through close discussions with our collaborators, we distilled a set of core design requirements that guided the development of ChartBlender.
  • R1
    Video-Suited Visualization Design. Visualizations should be designed to present data clearly while preserving the visual coherence of the scene, minimizing interference with the underlying video content.
  • R2
    Depth-Aware Embedding. Visualizations should be embedded into the video by aligning with camera depth and adopting a perspective consistent with the video scene. Such alignment ensures that the visualizations appear perceptually native to the scene, rather than as artificial post-production overlays.
  • R3
    Motion Synchronization. As the video plays, the visualizations should move in synchrony with the video, including the movement of both the camera and objects in the scene. To achieve this synchronization, it is necessary to track camera and object motion accurately.
  • R4
    Coordinated Multi-Chart Embedding. The system should support the integration of multiple visualizations within a single video, allowing authors to coordinate their spatial placement and temporal sequencing. This enables richer data storytelling and ensures that each chart appears at the appropriate moment and location in the visual narrative.
Figure 1: ChartBlender system consists of two modules: Chart Editor and Chart Embedder.
To fulfill these requirements, we developed the ChartBlender system. Fig. 1 illustrates the system architecture and the embedding pipeline. The system consists of two main modules: (1) the Chart Editor, and (2) the Chart Embedder. Generally, a user starts creating a visualization-embedded video by uploading a dataset and interactively configuring the visualization using the Chart Editor. This module allows users to customize various properties such as data mapping and style editing (R1). Afterward, users have a real-time preview of the chart’s 3d placement and appearance within the video to the Chart Embedder (R2). The Chart Embedder incorporates a tracking algorithm that synchronizes visualizations with motion in the video and operates without requiring any camera metadata. It first performs depth estimation on the uploaded video, then anchors the visualization to the scene based on the user’s selected tracking mode and target (R3). Finally, users can further adjust the chart, trim the timeline, and incorporate additional elements to produce smoother, more engaging, visualization-enhanced videos (R4). Collectively, these modules provide a lightweight yet powerful interactive workflow that automates chart embedding, reduces manual effort, and lowers the barrier for users. At the same time, the system supports the creation of visually coherent visualizations that integrate seamlessly into dynamic video scenes. While the Chart Editor functions as a relatively standard visualization creation module, in the following section, we focus on the Chart Embedder, which constitutes the core contribution of our system.
4 Chart Embedder
In this section, we present a tracking algorithm for synchronizing visualizations with video motion. The algorithm supports two scenarios: (1) keeping visualizations stably anchored to the background as the camera moves (camera tracking), and (2) allowing visualizations to follow moving objects within the scene (object tracking).
As shown in Fig. 2, given an input video consisting of $N$ frames and a user-specified target point $(x,y)$, our goal is to generate a sequence of 3D pose matrices $O_{t}$. Each pose matrix represents the position and rotation of the tracked point in 3D space at frame $t$. It begins with depth estimation (Fig. 2 (a)) from the input video frames, which produces per-frame depth maps. These maps are used in two parallel modules: camera tracking (Fig. 2 (b)), which computes frame-to-frame transformations and accumulates them into a global camera pose; and object tracking (Fig. 2 (c)), which tracks a user-specified point across frames and combines its 2D trajectory with depth information to estimate a smoothed 3D object pose. Finally, the estimated object pose is applied to the visualization, which is then overlaid onto the video scene.
Figure 2: The pipeline of the Chart Embedder.
4.1 Depth Estimation
Our system employs a monocular depth estimation model, Metric3D[38], to generate per-frame metric depth maps directly from original video frames. The major advantages of adopting Metric3D are twofold: (1) it predicts absolute metric depth without requiring ground-truth depth or camera intrinsics, enabling zero-shot deployment on in-the-wild videos; (2) it achieves high generalization capability across scenes, owing to its large-scale training on diverse data.
Given these benefits, we apply Metric3D to each of the $N$ video frames to generate a sequence of dense depth maps $D_{n}$, $n\in{0,1...,N-1}$. For a user-specified 2D point $(x,y)$, we can extract its first frame’s depth value from the depth map as $d=D_{0}(x,y)$. This depth information serves as a crucial foundation for the subsequent camera tracking and object tracking processes, enabling accurate spatial reasoning in both modules.
Figure 3: An illustration of the two motion embedding modes supported by our system: Camera Tracking and Object Tracking. In the Camera Tracking mode, the visualization remains anchored to a fixed point in the scene, attaching a chart to a stationary fence in the video. In contrast, Object Tracking binds the visualization to a moving object, attaching a chart to a galloping horse.
4.2 Camera Tracking
In ChartBlender, camera tracking ensures that embedded 3D visualizations remain spatially consistent with the static element throughout the video (Fig. 3(a)). To achieve this, we estimate the camera’s motion through a three-step process. First, we perform (1) relative pose estimation to compute the geometric transformation between adjacent video frames, capturing how the camera moves over time. This information is then used in (2) camera pose estimation, accumulating relative transformations to recover the camera’s trajectory in the scene. Finally, we conduct (3) 3D pose recovery to infer the camera’s position and orientation in 3D space, enabling visualizations to be accurately projected.
Relative pose estimation. To understand how the camera moves from frame to frame, we should estimate the relative transformation between each pair of consecutive video frames. These pairwise transformations form the foundation for reconstructing the full camera trajectory. We employ an RGBD odometry model provided by Open3D [39, 40] to compute the relative camera motion between adjacent RGBD frames. Given two consecutive frames $n$ and $n+1$, RGBD odometry estimates a transformation matrix. Specifically, we extract a transformation matrix $T_{n\rightarrow n+1}$ that aligns the RGB and depth images. Each transformation is a $4\times 4$ homogeneous transformation matrix composed of a rotation matrix $R_{n\rightarrow n+1}$ and a translation vector $t_{n\rightarrow n+1}$:
$T_{n\rightarrow n+1}=\left[\begin{array}[]{cc}R_{n\rightarrow n+1}&t_{n\rightarrow n+1}\\ 0&1\end{array}\right]$ (1)
Camera pose estimation. While relative poses describe local motion between frames, embedding visualizations in a stable, world-aligned coordinate system requires each frame’s absolute camera pose. This enables consistent rendering of the 3D object across the video sequence. We compute the absolute camera pose $M_{n}$ at each frame $n$ by chaining all previous relative transformations starting from the first frame:
$M_{n}=T_{1\rightarrow 2}T_{2\rightarrow 3}\cdots T_{{n-1}\rightarrow n}$ (2)
This accumulation step reconstructs the global camera trajectory, allowing us to express each frame’s viewpoint in the same coordinate system.
3D pose recovery. To render the 3D visualization as if it were physically anchored in the scene, we must compute its position in the global world coordinate system. Given a user-specified 2D point $(x,y)$ in the image and its corresponding predicted depth $d$, we first back-project the point into the 3D camera coordinate system using the inverse of the camera intrinsic matrix $K^{-1}$. The resulting 3D point in the camera frame is:
$P_{\text{n}}=K^{-1}\cdot\left(d_{0}\cdot\tilde{o}_{n}\right)$ (3)
where $P_{n}$ denotes the 3D position of a tracked point at frame $n$, $\tilde{o}_{n}=[x,y,1]^{T}$ is the homogeneous pixel coordinate. Next, we transform this point into the world coordinate system using the inverse of the absolute camera pose $M_{n}$, computed from accumulated relative poses. The object’s global pose $O_{n}$ is then given by:
$O_{n}=M_{n}^{-1}\cdot P_{\text{n}}$ (4)
This transformation places the 3D visualization at a consistent location in the world coordinate system, regardless of camera motion. It ensures that the visualization remains spatially stable and visually aligned with the static background in the video.
4.3 Object Tracking
In ChartBlender, the goal of object tracking is to enable visualizations to remain spatially aligned with moving elements in the video scene (Fig. 3(b)). To achieve this, we estimate the 3D motion of target objects from monocular video input through a three-stage pipeline. First, we perform (1) 2D point tracking to extract the object’s pixel-wise trajectory across frames, capturing how it moves within the image plane. This trajectory serves as the basis for the second stage, (2) 3D translation estimation, which lifts the 2D motion into 3D space to approximate the object’s position relative to the camera over time, allowing visualizations to maintain consistent placement as the object moves. Finally, we perform (3) 3D rotation estimation to recover the object’s orientation changes, ensuring that embedded visualizations can rotate coherently with the object’s local motion, further enhancing the perceptual integration.
2D point tracking. We begin by identifying salient points on dynamic objects and tracking their 2D positions across frames. To accomplish this, we use BootsTAP [41], a high-precision keypoint tracking algorithm that leverages dense visual correspondence estimation. For each target point $(x,y)$, the tracker generates a pixel-level trajectory over time, denoted as $p_{n}=(u_{n},v_{n})\rightarrow p_{n+1}=(u_{n+1},v_{n+1})$. The 2D trajectories capture the apparent motion of object features in image space and form the foundation for estimating real-world displacement in 3D, which is critical for subsequent motion reconstruction.
3D translation estimation. To reconstruct the spatial motion of tracked points, we lift the 2D trajectories to 3D space. For each frame, the 2D trajectory is projected into 3D space using the depth map of the tracked point in the camera coordinate system.
$d_{n}=D_{n}(p_{n})$ (5)
However, monocular depth estimates are inherently noisy and temporally unstable due to factors such as low-texture regions, occlusions, and variations in lighting. These inconsistencies can lead to jitter and discontinuities in the reconstructed 3D trajectories, especially when tracking fine-grained object motion. Such artifacts degrade both the physical plausibility and visual coherence of the motion path, undermining downstream tasks like 3D alignment and visualization anchoring. To address this, we apply temporal smoothing to the depth signal using quadratic regression over a sliding window:
$\hat{d_{n}}=\text{QuadraticRegression}(d_{n})$ (6)
As previously defined in Equation (3), we recover the 3D translation vector $t_{n}$ of the target point in the camera coordinate system by back-projecting its 2D image coordinate using the estimated depth and the intrinsic matrix K.
3D rotation estimation. Finally, we estimate object rotation based on the direction of motion in 3D space to ensure that visualizations remain properly oriented when attached to moving objects. Specifically, we define the per-frame motion vector as $V_{n}=P_{n+1}-P_{n}$. We obtain the rotation vector $R_{n}$ by constraining the object’s local X axis to follow $V_{n}$. To support stable rotation estimation, we first regularize the raw 3D point trajectories by minimizing sudden changes in displacement. This is implemented as a second-order temporal smoothing loss:
$\min_{\hat{P}}\sum_{n}\|\hat{P}_{n+1}-\hat{P}_{n}\|^{2}+\lambda\sum_{n}\|\hat{P}_{n+2}-2\hat{P}_{n+1}+\hat{P}_{n}\|^{2}$ (7)
Next, we apply the same principle in the velocity domain to ensure that the direction of motion also varies smoothly over time. The corresponding velocity regularization is formulated as:
$\mathcal{L}(V_{n})=\sum_{n}\|V_{n+1}-V_{n}\|^{2}+\lambda\sum_{t}\|V_{n+1}-2V_{n}+V_{n-1}\|^{2}$ (8)
This second-order smoothing constraint encourages both consistent motion direction and smooth acceleration profiles, mitigating jitter and improving temporal coherence in the estimated trajectories. Given the smoothed translation vector and rotation vector, the final object pose at each frame is defined as $O_{n}=\{R_{n}\mid t_{n}\}$.
5 ChartBlender
Figure 4: The interface consists of three main components: (1) the Chart View for creating and editing visualizations, (2) the Video View for embedding charts and previewing the embedding result, and (3) the Timeline View for adding, removing, and trimming chart segments.
In this section, we introduce the interface design, visualization design, and interaction.
5.1 User Interface
The user interface, as shown in Fig. 4, consists of three major views: the Chart View (Fig. 4-1), the Video View (Fig. 4-2), and the Timeline View (Fig. 4-3). In the Chart View, users can upload a dataset, after which the raw data is displayed directly within the view. Users can then choose a chart type (e.g., bar chart, line chart, area chart, and text) and define encoding mappings by dragging data fields to visual channels such as the x-axis, y-axis, timestamp, or title. To refine the visual presentation, users can adjust stylistic properties including line thickness, color, and theme (R1). The Video View provides a real-time preview of the current video frame. Users can manipulate the visualization’s position, rotation, and scale in 3D space, enabling precise spatial embedding within the scenes (R2). To synchronize visualizations with video motion, users first select a motion mode, either camera tracking or object tracking (Fig.4-2(a)), and then specify target points in the video as tracking anchors (Fig.4-2(b)). After configuration, clicking the Render button (Fig.4-2(c)) triggers the system to compute the visualization’s motion path and composite it into the video (R3). Once rendering is complete, users can click the Play button to preview the embedded visualization in motion, ensuring alignment and visual consistency.
The Timeline View serves as an editing panel that supports temporal coordination and layer management. On the left side of the interface, users can add, delete, or lock chart layers to organize content across multiple visualization tracks. On the right side, a Timeline allows users to control the duration and timing of each chart via intuitive drag, stretch, and trim interactions (R4). Finally, by clicking the Export button, users can generate the composite video with the embedded visualizations synchronized to the video.
5.2 Visualization Design
Our goal is to design visualizations that are well-suited for integration within videos. This requires rethinking how charts are structured and styled to maintain both visual clarity and contextual harmony in videos[42]. To this end, we collected a total of 200 high-quality videos and manually segmented them into 502 clips, each containing at least one embedded visualization. As shown in Fig. 6, these videos span ten diverse domains. We also analyzed the most commonly used chart types across the dataset. Furthermore, we identified several recurring design patterns in how visualizations are embedded within real-world video content. These patterns are summarized as follows.
Figure 5: Design space for visualizations suitable for embedding in video, which illustrates five representative visualization forms combined with two motion anchoring strategies (object tracking and camera tracking). It reveals common design patterns for how visualizations are structured and integrated within dynamic video scenes across diverse real-world applications.
Figure 6: Distribution of chart types and video domains.
First, a majority of charts (130 cases) incorporate background to improve readability, with a subset (21) employing semi-transparent backgrounds to maintain visual access to the underlying scene. Notably, gridlines were selectively applied to support data interpretation, with horizontal guides appearing more frequently (109) than vertical ones (43), while 64 charts omitted them entirely. Second, very few charts (only 4) included legends, indicating a general preference for minimalist presentations in time-constrained video contexts. Nevertheless, essential structural components such as axes (retained in 171 charts) and titles (present in 168 charts) were commonly preserved. Third, when visualizations were associated with specific objects in the scene, designers frequently employed visual linking strategies to establish a clear connection between the chart and its target (68). These strategies included directional cues such as arrows, lines, and bounding boxes. The choice of pointer type often varied depending on scene complexity: in simple scenes with few moving objects, minimal indicators such as triangles were typically sufficient; in more complex scenes with multiple overlapping motions, designers commonly used bounding boxes or anchored pointers to more clearly distinguish the tracked target.
These findings highlight a set of implicit yet widely adopted design patterns for video-suited visualizations design, which we synthesize into a structured design space, as illustrated in Fig. 5. This design space is organized along two key dimensions: the motion mode and the visualization type. It distinguishes between two modes of motion (object tracking and camera tracking) corresponding to whether the visualization follows a moving target in the scene or remains fixed relative to the global environment. Horizontally, it categorizes common visualization types, including text annotations, line charts, area charts, bar charts, and 3D charts, as well as their corresponding linking and layout strategies (e.g., aiming lines, bounding boxes, or indicators). Based on this design space, we developed a visualization chart library using Three.js, providing a collection of chart templates optimized for embedding in dynamic video scenes.
5.3 Interaction
To support efficient visualization authoring and seamless synchronization within video, we propose a set of interactions in ChartBlender:
Upload. Users can upload datasets and video files either by clicking the upload button or by dragging and dropping files into the designated panel.
Data Mapping. Data attributes can be mapped to visual channels through direct manipulation: users drag data fields onto chart elements to define encoding relationships.
Zoom and Pan. Users can interactively adjust the visualization view: left-click and drag to reposition charts, scroll to zoom in and out, and right-click and drag to rotate.
Chart and Style Switching. Users can switch between different chart types via tabs in the interface. They can also adjust visual styling, such as color schemes, by clicking style options or using a color picker tool.
Anchor Selection. To ensure accurate chart placement, users can click on static or dynamic objects in the video to define visual anchor points. These anchors are used to bind visualizations to specific elements in the scene, supporting consistent spatial alignment across frames.
Track Editing. Users can manage chart tracks via the timeline view. Tracks can be added or removed by clicking the corresponding buttons in the toolbar on the left. On the right, users can adjust the timing and duration of each chart by dragging chart tracks directly along the timeline.
6 Evaluation
We evaluated the ChartBlender and the key techniques through three approaches: (1) a quantitative experiment to assess the precision of camera tracking, (2) a quantitative experiment to assess the precision of object tracking, and (3) domain expert interviews to evaluate the overall usability of the system.
6.1 Camera Tracking Precision
In our system, camera tracking is implemented using RGBD odometry, which estimates frame-by-frame camera motion by leveraging both RGB appearance and depth information. As a result, the quality of depth estimation plays a critical role in the precision of camera tracking. To evaluate this dependency, we compare how different monocular depth models affect the overall pose estimation performance of the RGBD odometry pipeline.
Experimental Setup. To evaluate the impact of depth estimation quality on camera pose recovery, we conducted a comparative experiment using different monocular depth models combined with RGBD odometry. Specifically, we compared Metric3D, which emphasizes metric depth precision, with ZoeDepth, which focuses on generalization and temporal consistency. We refer to the two pipelines as Metric3D + RGBD odometry (our proposed method) and ZoeDepth + RGBD odometry (baseline). We selected a subset of the DriveTrack dataset [43], consisting of 18 video sequences (1,806 frames in total), where camera motion was sufficiently pronounced and affected by dynamic foreground objects. For each method, predicted depth maps were fed into RGBD odometry implemented in Open3D [39], which uses a hybrid geometric-photometric optimization [40] to estimate per-frame camera extrinsics. A fixed intrinsic matrix of the form [1000, 1000, width/2, height/2] was used for all sequences.
Evaluation Metric. To quantitatively evaluate the precision of our camera tracking approach and assess the influence of different depth estimation models, we measured the geometric discrepancy between the estimated camera poses and the ground-truth extrinsics provided by the TAPVid-3D dataset [34]. For each frame, we compared the predicted camera pose $(R_{\text{pred}},t_{\text{pred}})$ with the corresponding ground-truth pose $(R_{\text{gt}},t_{\text{gt}})$, and computed both rotation and translation errors. The rotation error was defined as the angular difference between the predicted and ground-truth orientations. Specifically, we first computed the relative rotation matrix $\Delta R=R_{\text{gt}}^{-1}R_{\text{pred}}$, and then calculated the rotation error as the angle of this rotation:
$\theta=\arccos\left(\frac{\operatorname{trace}(\Delta R)-1}{2}\right),$ (9)
This formulation captures the minimum angular difference between the two rotation matrices in 3D space. The translation error was defined as the Euclidean distance between the predicted and ground-truth camera positions:
[…]
Together, these two metrics quantify both orientation and positional discrepancies in the estimated camera trajectories. To evaluate overall precision and robustness to outliers, we report both the mean and median rotation and translation errors computed across all frames in the dataset.
Results and Analysis. As shown in Table I, Metric3D consistently outperformed ZoeDepth in both rotation and translation precision when paired with RGBD odometry, achieving lower mean and median errors across all frames. Although both methods operate under the same fixed intrinsic setting, Metric3D provided more reliable metric depth, enabling more accurate frame-by-frame camera pose estimation. Another key reason for adopting Metric3D is its robustness in unconstrained environments where intrinsic calibration is unavailable, a common challenge in user-generated or in-the-wild videos. Despite the use of an approximate intrinsic matrix, the Metric3D + RGBD odometry pipeline maintained strong performance.
TABLE I: Comparison on camera tracking. (Camera pose estimation errors)
Method Rotation Error (degrees) Translation Error (m)
Mean ↓ Median ↓ Mean ↓ Median ↓
Metric3D + RGBD odometry (ours) 8.641 8.851 6.518 6.432
ZoeDepth + RGBD odometry 10.544 11.061 7.598 7.610
6.2 Object Tracking Precision
We evaluated our system’s object tracking techniques by assessing the precision of the reconstructed 3D motion trajectories for dynamic objects. This reflects how well the system can estimate object movement in physical space.
Experimental Setup. In our system, object tracking is implemented by combining 2D keypoint tracking with monocular depth estimation to recover the 3D trajectories of dynamic objects across time. This process is essential for enabling visualization alignment with moving targets in the scene. The overall tracking precision depends on the quality of both the 2D keypoint trajectories and the estimated depth maps used for lifting to 3D. To evaluate this component, we conducted a systematic comparison of different tracking configurations by varying the underlying 2D tracking and depth estimation models. Specifically, we considered two keypoint tracking models, TAPIR [44] and BootsTAPIR [41], and two depth estimation models, ZoeDepth and Metric3D, resulting in four combinations in total.
We performed this evaluation on a subset of the PStudio dataset [45], consisting of 36 video sequences (10,800 frames in total). These sequences feature clearly observable dynamic objects with prominent motion and minimal camera movement, allowing us to isolate object trajectory estimation from background motion. For each combination, 2D keypoints were first tracked across frames using the selected tracking model. These trajectories were then lifted into 3D space by back-projecting the 2D points with the corresponding depth estimates.
Evaluation Metric. We adopted the Average 3D Point Distance (APD3D) metric, introduced in TAPVid-3D [34], to evaluate the precision of 3D point tracking. This metric quantifies the proportion of predicted 3D points that lie within a specified Euclidean distance threshold of their corresponding ground-truth positions, averaged over time and across tracked points. A higher value reflects better spatial precision and temporal consistency of the reconstructed trajectories.
Formally, the metric is defined as:
$\mathrm{APD}_{3D}\equiv\frac{1}{V}\sum_{i,t}v_{t}^{i}\cdot\mathbf{1}\left(\lVert\hat{P}_{t}^{i}-P_{t}^{i}\rVert<\delta_{3D}(P_{t}^{i})\right),$ (11)
where $\hat{P}_{t}^{i}$ and $P_{t}^{i}$ denote the predicted and ground-truth 3D positions of the $i$-th point at time $t$, $v_{t}^{i}$ is a visibility indicator, and $V$ is the total number of visible points. To quantify performance across different levels of precision, we report APD3D under a fixed set of distance thresholds: $[0.04,0.08,0.16,0.32,0.64]$.
Results and Analysis. As shown in TABLE II, BootsTAPIR consistently outperformed TAPIR across all configurations, confirming its superior capability in 2D temporal keypoint tracking. With regard to depth estimation, our primary objective is not to reconstruct metrically accurate 3D geometry, but to generate perspective-consistent visualizations that preserve spatial coherence between foreground and background elements. In this context, moderate depth errors are acceptable, provided that the overall depth structure remains stable and visually plausible.
[…]
BootsTAPIR + Metric3D (ours) 0.3424
TAPIR + Metric3D 0.3416
BootsTAPIR + ZoeDepth 0.3609
TAPIR + ZoeDepth 0.3595
Figure 7: Five example videos with embedded visualizations generated using ChartBlender
6.3 Expert Interview
To further evaluate the usability of ChartBlender, a series of interviews with five data journalists (denoted as E1-E5) were performed. They had over 5 years’ video editing experiences on average, and were very familiar with chart-embedded videos. All participants provided informed consent prior to the study.
Procedure. The interviews were performed via an online meeting system. Each interview started by a 10-minute introduction about the system. Participants were first introduced to our system through a brief tutorial delivered by authors, and were given two task:
Task 1 - Goal-oriented (10 minutes). Participants are shown an example video that they have to reproduce. The result should consist of a bar chart that remains fixed in the field and an area chart that moves along with the vehicle.
Task 2 - Open-ended (30 minutes). Participants are asked to create two novel data videos by choosing among 10 short video clips (10” duration on average).
Finally, the experts were asked to finish a questionnaire, followed by an interview after they explored the functionalities of ChartBlender. The whole process lasted for about one hour, and the interviews were recorded for analysis.
Result. Fig. 7 presents sample videos generated by expert users using our system. In follow-up interviews, the experts offered a range of insightful comments, which are summarized as follows:
Fast Crafting. All experts agreed that our system was highly effective in supporting both camera tracking and object tracking. In particular, E1 commented that “This system can accomplish the work of several hours, or even days, in just two or three minutes.” E2 mentioned that “This system allows me to see the real-time effect of chart edits in the video… I no longer need to repeatedly switch between project files, which used to waste a lot of time.” E5 also noted that “The speed of result generation is so fast that I can quickly create multiple demos to observe where the chart embedding works best in the video.”
Accurate Results. Most experts were very satisfied with our results, both for camera and object tracking (E1, E2, E4, E5). E1 commented, “The results are directly usable.” E2 suggested that “The object tracking techhnique amazed me; it not only takes the camera pose into account but also integrates the movement of the objects.” E5 mentioned, “This result is more accurate than manual alignment.”
A Professional and Elegant System. All the experts were satisfied with the system. Additionally, all of them mentioned that the system was very user-friendly and that they would recommend it to others. Specifically, they appreciated how seamlessly the system integrates chart creation and embedding into videos, accomplishing highly professional tasks (E1, 2, 5). E1 mentioned that “This system can complete highly professional tasks without requiring extensive skills or knowledge.” E5 noted that “the chart can be translated and rotated in 3D, and it integrates seamlessly with the video.” They were also satisfied with the chart creation aspect. E4 noted that “the data mapping feature eliminated the need for selecting data and creating visualizations in 3D software.” E5 commented that “The chart creation process is very user-friendly for me. There are many expressive templates available for me to use … Previously, I had to spend a lot of time on creating visualizations.”
7 LIMITATIONS AND FUTURE WORK
Despite the encouraging evaluation results, several discussions and limitations were identified during the design and implementation process or mentioned by the experts during interviews. We hope to guide potential future research directions by pointing out these limitations.
Context-Aware Visualization Generation. Several experts (E1, E2, E4) expressed the need for a more intelligent system capable of understanding both dataset and video semantics. For instance, the system could automatically suggest visualization color schemes aligned with the video or incorporate domain-specific visual metaphors, such as rendering bars in a bar chart as crops in agricultural contexts. It could also generate appropriate chart titles based on the video narrative and the underlying data. Future work may explore fine-tuning large language models (LLMs) to support automatic topic analysis, enabling the system to generate context-aware visualizations with optimized design elements, enhancing the coherence and quality of data-driven videos.
Layered Rendering. Although the current system supports depth-aware relationships between charts, Several experts (E1, E3) highlighted the potential of masking functionality that enables visualizations to be occluded by foreground objects in the video. Such a feature would allow charts to appear partially or fully behind scene elements, enhancing realism and spatial coherence.
Tracking Editing. Currently, the system automatically generates motion trajectories for embedded charts based on anchor points. A promising extension is to introduce a visual trajectory editor that renders the underlying coordinate array as editable curves within a canvas or SVG-based interface. Leveraging techniques such as spline interpolation, draggable control points, and direct path manipulation, users could intuitively modify the motion path. Such an interface would offer frame-level control over trajectory shapes, allowing insertion, deletion, and adjustment of keypoints. This would significantly enhance the flexibility and expressiveness of the authoring workflow.
8 Conclusion
In this paper, we introduced ChartBlender, an integrated system for authoring, embedding, and synchronizing data visualizations within dynamic video content. By combining computer vision algorithms for camera and object tracking, we enable synchronization between visualizations and video motion. In addition, the system incorporates an interactive workflow to support easy and efficient authoring of videos with visualizations embedded. Quantitative evaluations demonstrate the effectiveness of our tracking algorithms in recovering spatially consistent motion trajectories. Additionally, interviews with domain experts confirm the system’s usability, expressiveness, and potential for professional media production. The evaluation showed the power of ChartBlender system and revealed several limitations of the current system, which will be addressed in the future.
ACKNOWLEDGMENTS
Nan Cao is the corresponding author. This work was supported by the National Key Research and Development Program of China (2023YFB3107100). We would like to thank all the reviewers for their valuable feedback.

### v2

@@
 Abstract
-Embedding data visualizations in video can enhance the communication of complex information.
-However, this process is often labor-intensive, requiring designers to adjust visualizations frame by frame manually.
-In this work, we present ChartBlender, a novel system that streamlines this process by enabling users to create data visualizations, embed them seamlessly into video scenes, and automatically synchronize them with both camera motion and moving objects.
-Particularly, ChartBlender incorporates a tracking algorithm that supports both object and camera tracking, ensuring robust alignment of visualizations with dynamic video content.
+Embedded data visualizations have emerged as a powerful narrative medium for conveying complex information within video footage.
+However, creating such content remains labor-intensive, as existing workflows rely on manual frame-by-frame adjustments to ensure spatial and temporal consistency.
+To address these challenges, we present ChartBlender, an interactive authoring system designed to streamline the creation, embedding, and automatic synchronization of data visualizations within video scenes.
+We develop a tracking pipeline that supports both object and camera tracking, ensuring robust alignment of visualizations with dynamic video content.
 To maintain visual clarity and aesthetic coherence, we also explore the design space of video-suited visualizations and develop a library of customizable templates optimized for video embedding.
-We evaluate ChartBlender through two controlled experiments and expert interviews with five domain experts.
+We evaluated ChartBlender through two controlled experiments and expert interviews with five domain experts.
 Results show that our system enables accurate synchronization and accelerates the production of data-driven videos.
-Index Terms:
-data visualization, video editing, object tracking, camera tracking, interface, visual storytelling.
-publicationid: pubid: $0$–$00$ 2021 IEEE
+publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
 1 Introduction
-Embedding data visualizations into videos has emerged as a powerful method for conveying complex information [1, 2, 3, 4, 5].
-Such a form of representation is increasingly used across various domains (e.g., sports, traffic, and journalism), allowing audiences to access supplementary information without being distracted from the primary video content[6].
-Despite its potential, the process remains labor-intensive: achieving spatial and temporal synchronization between visualizations and video typically requires manual, frame-by-frame alignment.
-Even a 10-second visualization-embedded video may demand up to a full day of manual work from designers.
-Motivated by a collaborative project with a major Chinese media organization to create a chart-embedded video, we conducted a comprehensive survey of existing tools and systems.
-On one end of the spectrum are professional tools like Adobe After Effects [7] and Blender[8].
-While powerful, these tools rely on manual parameter tuning and lack intelligent assistance, making them difficult to use efficiently.
-On the other end, recent advances in Augmented Reality (AR) and Virtual Reality (VR) have inspired research systems that allow casual users to overlay visualizations on videos for sharing quickly [9].
-These systems prioritize ease of use and rapid output but are designed primarily for end users with simple needs.
-As a result, they fall short in supporting high-quality, professionally produced content.
-Specifically, they lack robust motion tracking for objects with diverse movement patterns, as well as precise control over visualization placement and editing capabilities essential for content creators and news organizations.
-Therefore, there is a pressing need for an interactive authoring system supports flexible, efficient embedding of data visualizations into complex video scenes and ensures accurate motion synchronization throughout the video.
-However, to design such an effective system, three challenges need to be tackled: (1) track the motion of videos accounting for both moving objects in the scene and camera pose variations, to ensure consistent synchronization between visualizations and the video scene; (2) design generalized visualizations to support clear and informative presentation of data within dynamic video contexts; and (3) enable user-driven authoring and refinement of visualizations through simple and integrated interaction.
-Addressing these challenges is essential for the precise and seamless embedding of data visualizations into real-world videos.
-To address these challenges, we introduce ChartBlender, a novel authoring system for embedding and automatically synchronizing data visualizations with video during playback.
-In particular, the user simply selects a target tracking point and a preferred tracking mode, and the system automatically generates the complete motion trajectory for visualization throughout the Users can edit the visualization at any time, including both its visual style and its 3D pose within the video scene.
+Embedding data visualizations into videos has emerged as a powerful way to communicate complex information [idvxNVSurvey, amini2015understanding, shi2021autoclips, 10891192, 7539328].
+This approach is now widely adopted in domains such as sports [iball], traffic [hsieh2016traffic], and journalism [covid19_youtube_2021].
+By integrating charts into footage, it can provide timely and contextual insights while allowing viewers to stay focused on the underlying footage [iball].
+Real-world examples of this visual medium can be seen in a video from The Economist [^1] https://www.youtube.com/watch?v=pFaGbtDJFwA&list=WL&index=8(4’35’’--5’20’’) (Fig.
+1(a)).
+In this segment, a line chart is anchored to a wall and remains synchronized with camera motion to illustrate behavioral changes during and after the pandemic.
+Figure 1: Real-world examples of data visualizations embedded in videos.
+(a) A frame from The Economist, where a line chart illustrates behavioral changes during and after the pandemic.
+(b) A frame from a national media organization in China, where a bar chart shows year-by-year growth in participant numbers.
+Despite growing interest, producing videos with embedded visualizations remains a challenging task.
+Our collaboration with a national Chinese media organization  [^2] http://edu.people.com.cn/n1/2023/0915/c1006-40078200.html(2’44’’,3’25’’) (Fig.
+1 (b)) revealed critical workflow bottlenecks.
+During the collaboration, we were tasked with embedding a 3D bar chart into an aerial landscape shot to visualize participant growth.
+The chart needed to remain synchronized with camera motion.
+However, to the best of our knowledge, no dedicated authoring tool provides motion-aware synchronization tailored to embedded data visualizations.
+In practice, we relied on general-purpose video-edit software and laborious manual keyframing of the chart’s pose and trajectory.
+Even minor revisions often triggered substantial rework, and producing a 10-second clip could take nearly a full workday.
+These bottlenecks directly informed our design requirements and system goals, highlighting an urgent need for an integrated, motion-aware authoring system for professional data-driven video production.
+Within the visualization community, interest in video-embedded visualizations is also increasing.
+However, existing tools do not meet the requirements of motion-aware visualization authoring in real-world production.
+For example, sports annotation systems [7539370, iball, swimflow] demonstrate the value of integrating tracked graphics into video narratives, yet they often rely on domain-specific pipelines, fixed visual encodings, or approximate 2D overlays rather than full scene-consistent integration.
+Meanwhile, some AR/VR-based authoring tools [8611113, VisTellAR] enable end users to quickly place charts onto video, emphasizing ease of interaction and shareability; but they offer limited visual and stylistic control, make precise alignment and post-capture refinement difficult, and scale poorly to multi-chart compositions.
+Commercial video editing platforms (e.g., After Effects [adobe_ae], Blender [blender], DaVinci Resolve [davinci]) provide motion tracking and compositing capabilities, but visualization embedding typically requires labor-intensive workflows, substantial expertise, and disjoint processes that separate data design from in-context visual authoring.
+Overall, while motion tracking is widely available, existing tools rarely couple it tightly with visualization authoring, limiting dynamic, context-aware chart placement in complex scenes with editable visual encodings.
+Therefore, there is a pressing need for an interactive authoring system that supports flexible and efficient embedding of data visualizations into complex video scenes while maintaining accurate motion synchronization throughout the video.
+However, to design such an effective system, three challenges need to be tackled: (1) robustly tracking motion by accounting for both object movement and camera pose variation to ensure consistent synchronization between the visualization and the scene; (2) supporting generalized chart designs that remain legible and informative under dynamic video conditions; and (3) enabling simple, integrated interactions for user-driven authoring and iterative refinement.
+Addressing these challenges is essential for precise and seamless visualization embedding in real-world videos.
+To address these challenges, we present ChartBlender, an authoring system that embeds and synchronizes data visualizations with video content.
+Users specify a tracking point and choose a tracking mode, and the system automatically generates the motion trajectory.
+At any time, users can adjust both the chart’s visual style and its 3D position in the scene.
 In this way, the system supports the human-machine collaboration to generate data-driven videos rapidly while ensuring the visualization appears visually coherent and seamlessly embedded into the video.
-The major contributions of the paper are as follows:
+The major contributions of the paper are as follows.
   • System.
@@
 It features an interactive interface that supports the end-to-end creation of visualization-embedded videos through an integrated authoring workflow.
+We also developed an algorithm pipeline to support robust visualization embedding for the system.
   • Design Space.
@@
 Based on this analysis, we develop a library of generalized visualization templates designed to convey complex data clearly and effectively within dynamic video contexts.
-  • Algorithms.
-We develop two algorithms to support robust visualization embedding in dynamic video scenes.
-The first estimates frame-by-frame camera poses, enabling visualizations to remain stably anchored to the background.
-The second computes and refines object trajectories across frames, allowing visualizations to accurately follow moving targets within the video.
+  • Evaluation.
+We demonstrate the utility of ChartBlender through a set of example videos created from real-world data, two controlled experiments that validate the tracking pipeline, and a series of interviews with five expert users.
 2 Related Work
-In this section, we survey the papers that are most relevant to our work, including professional video editing software, embedded visualization, and point tracking.
+In this section, we review the recent studies that are most relevant to our work, including embedding visualizations in video, video editing software, and point tracking.
+2.1 Embedding Visualizations in Video
+Embedding visualization, as a form of embedded data representation, places data-encoded graphics in close spatial correspondence with real-world referents (e.g., objects or locations), enabling viewers to interpret data in context rather than consulting a separate chart view [7539328].
+When the target medium is video, this coupling should be preserved under continuous camera and scene motion, making the problem a form of visualization in motion[yao2020].
+In this setting, relative motion and limited viewing time can affect how accurately and efficiently viewers read values, while professional production further requires stable spatiotemporal alignment and reliable legibility under changing viewpoints, cluttered backgrounds, and occlusions.
+Early efforts such as SmartShot [smartshot] examined static embedding by optimizing chart placement to reduce occlusion of salient content and to balance aesthetic integration with readability.
+Beyond appearance, however, realistic embedding in dynamic 3D footage requires viewpoint-stable registration.
+As a result, perspective consistency, correct depth ordering, and occlusion handling become essential.
+To preserve visual coherence, embedded charts should follow camera trajectories and object dynamics; accordingly, a range of techniques and systems have been proposed [7539370, 10.1145/3411764.3445337, 10.1145/3411764.3445344, 10.1109/TVCG.2021.3114774].
+For example, SwimFlow [swimflow] embeds animated visualizations in swimming videos by annotating key positions and interpolating trajectories, which works well for predictable motion but is less general for arbitrary scenes.
+iBall [iball] overlays dynamic cues in basketball videos and adapts displays based on gaze; however, under camera motion the graphics are only approximately anchored and do not provide perspective-consistent 3D registration.
+Overall, many existing systems rely primarily on 2D overlays and tracking, offer limited depth-aware occlusion reasoning, and are tailored to specific domains.
+More recent work has explored general-purpose authoring systems that synchronize charts with arbitrary videos.
+With advances in AR/VR, it has become easier to create motion-following overlays.
+Chen et al.
+[8611113] introduced MARVisT, a mobile authoring tool that lets non-experts bind data to virtual glyphs and real objects, and systems such as VisTellAR [VisTellAR] enable casual users to overlay visualizations on videos quickly.
+However, these end-user tools typically trade expressive control and production robustness for accessibility: they provide limited chart grammars and styling controls, make precise alignment and post-capture refinement difficult, and scale poorly to multi-chart compositions.
+In this work, we introduce ChartBlender as an interactive authoring system for embedding and synchronizing charts in video.
+ChartBlender supports 3D object tracking and perspective-consistent registration to anchor charts stably in dynamic scenes.
+Unlike prior systems that rely on fixed overlays or narrow workflows, ChartBlender enables flexible, in-context authoring: designers can edit chart types, styles, and data mappings directly within the video while maintaining precise control over placement and timing.
+This integration streamlines the creation of motion-aware embedded visualizations for professional production.
 2.2 Video Editing Software
-Video editing is a broad field that involves various techniques, technologies, and tools used to create, manipulate, and refine video content, which often pose barriers to users [10, 11].
-Professional video editing software offers powerful tools to insert animated objects into videos.
-Traditionally, users must manually keyframe an object’s position, rotation, and scale over time, with smooth transitions achieved through interpolation.
-While recent advances have introduced automated tracking methods, such as 3D camera estimation and object tracking, these techniques, available in tools like Adobe After Effects, Blender, and Nuke [7, 8, 12], still require a level of expertise that can be challenging for beginners.
-At the same time, the growing demand for data-driven videos highlights a gap in current workflows: mainstream video editing tools lack integrated support for data visualization.
-Users often switch between separate tools or project files to create visualizations, making it difficult to assess whether a chart fits naturally within a given video context.
-Currently, no tool allows for direct editing of chart styles within the video environment while supporting real-time assessment, motion synchronization, and object tracking.
-This creates a pressing need for authoring tools that bridge the gap between visualization and video editing.
-2.2 Embedded Visualization in Video
-Embedded data visualization in videos has become increasingly popular [13, 14, 15, 16].
-However, this process often requires users to manually create keyframes for positioning the visualizations within the video.
-Recently, there has been increasing research on intelligent systems that can help embed visualizations in videos [17, 18, 6, 19].
-Early works primarily focused on embedding static charts into videos.
-For instance, SmartShot [17] developed a system that optimizes the placement of charts to avoid obstructing the visual focus, ensuring aesthetic balance and readability.
-SwimFlow [18] explored how to embed visualizations into swimming competition videos.
-Their system allows charts to animate by using manual annotations to mark athletes’ starting positions and interpolating to calculate their motion trajectories.
-However, this approach is clearly limited when dealing with more complex motion patterns.
-Recent advancements in computer vision now enable the embedding of visualizations directly into sports videos, offering a more dynamic and adaptive solution.
-For instance, iBall [6] embeds dynamic visual data cues into basketball videos and automatically adjusts the information display based on the user’s gaze.
-However, when the camera moves, the charts can only approximate the motion rather than being stably anchored to the scene.
-Additionally, these methods typically involve 2D overlays and tracking.
-In contrast to these domain-specific solutions, our approach tracks 3D trajectories of arbitrary objects and supports 3D effects, including perspective transformations and occlusions of the inserted visualizations.
-Moreover, existing systems often lack the flexibility to freely create and edit visualizations to meet user-specific needs.
-In our system, users can directly edit the types, styles, and data mappings of charts on the video, streamlining the creation process and offering greater customization.
+Nowadays, modern video editing and VFX tools provide motion tracking capabilities that allow designers to attach graphics or objects to dynamic scenes.
+After Effects [adobe_ae], for example, includes point tracking, spline-based tracking, and 3D camera tracking, enabling powerful compositing; in practice, however, achieving stable, perspective-consistent overlays, especially in the presence of fast motion or occlusions, often requires substantial manual refinement.
+Blender [blender] provides camera tracking and 3D scene integration within a full 3D authoring environment, but producing polished results typically requires expertise and nontrivial setup.
+Final Cut Pro [final_cut] offers an accessible, machine-learning-based object tracker for effects such as blurring or target-following, which is efficient for lightweight editing but provides less fine-grained control.
+DaVinci Resolve’s Fusion module [davinci] supports planar tracking and 3D camera tracking in a node-based workflow, yet its learning curve and context switching between editing and compositing can hinder iterative authoring.
+Foundry Nuke [nuke] is an industry-standard compositor with advanced planar tracking and 3D camera solving; it offers high precision but is costly and complex, and is therefore most commonly used in professional studio pipelines.
+Across these tools, automatic tracking alone often falls short for chart-embedded video authoring because (1) tracking is not visualization-specific and frequently demands frame-by-frame refinement for chart stability, (2) professional workflows impose steep learning curves and high authoring overhead, and (3) chart design and tracking are decoupled, making in-context preview and iterative refinement of visualization appearance cumbersome.
+By contrast, ChartBlender offers a lightweight and learnable workflow with interactive controls that allow authors to adjust chart styling and placement while immediately previewing the results in context.
 2.3 Point Tracking
-Tracking arbitrary points in videos has witnessed substantial progress in recent years [20, 21, 22], with methods such as TAP-Net [23] demonstrating the ability to reliably follow both static and dynamic points throughout a video sequence.
+Tracking arbitrary points in videos has witnessed substantial progress in recent years [zhou2020tracking, gauglitz2011evaluation, 10.1007/978-3-030-58548-8_28], with methods such as TAP-Net [hu2020tap] demonstrating the ability to reliably follow both static and dynamic points throughout a video sequence.
 These approaches typically leverage appearance-based cues and temporal consistency to maintain accurate correspondences, even under significant motion and occlusion.
@@
 To achieve this, it becomes necessary to disentangle camera motion from scene dynamics, which calls for robust camera pose estimation.
-Fortunately, a variety of mature techniques now exist for estimating camera pose from monocular or multi-view video streams [24, 25], ranging from classical SLAM-based pipelines [26, 27, 28] to deep learning methods[29, 30, 31, 32, 33].
+Fortunately, a variety of mature techniques now exist for estimating camera pose from monocular or multi-view video streams [xu2024critical, guan2024survey], ranging from classical SLAM-based pipelines [campos2021orb, vidal2018ultimate, 7219438] to deep learning methods[shavit2019introduction, schonberger2016structure, Hidalgo-Carrio_2022_CVPR, zhou2017unsupervised, en2018rpnet].
 The integration of 2D point tracking with camera pose estimation enables a more unified paradigm: tracking arbitrary points in 3D space.
-This formulation has recently attracted increasing attention [34, 35, 36], as researchers extend traditional 2D tracking frameworks into the 3D domain.
-For instance, SpatialTracker[36] is the first work specifically designed for the task of 3D point tracking.
+This formulation has recently attracted increasing attention [koppula2024tapvid, yu2023videodoodles, xiao2024spatialtracker], as researchers extend traditional 2D tracking frameworks into the 3D domain.
+For instance, SpatialTracker[xiao2024spatialtracker] is the first work specifically designed for the task of 3D point tracking.
 It estimates the 3D trajectory of arbitrary pixels by combining image features with monocular depth estimation, and propagates information through a graph neural network (GNN) to model long-range spatiotemporal motion.
-Similarly, TAPIR-3D[34] is an experimental extension of the original 2D point tracking model TAPIR, designed to output 3D point trajectories.
-It initializes keypoints in each frame, refines their 2D tracks using temporal context from adjacent frames, and subsequently lifts them to 3D space using depth estimation models such as ZoeDepth [37].
-We propose a novel approach that enhances the accuracy of 3D point tracking by integrating it with the practical requirements of embedding visualizations into real-world video content.
+Similarly, TAPIR-3D[koppula2024tapvid] is an experimental extension of the original 2D point tracking model TAPIR, designed to output 3D point trajectories.
+It initializes keypoints in each frame, refines their 2D tracks using temporal context from adjacent frames, and subsequently lifts them to 3D space using depth estimation models such as ZoeDepth [bhat2023zoedepth].
+We propose a pipeline that enhances the accuracy of 3D point tracking by integrating it with the practical requirements of embedding visualizations into real-world video content.
 Our system enables users to anchor visualizations to fixed locations within the scene or dynamically attach them to moving objects, regardless of camera motion or scene complexity.
@@
 3 System Design and Overview
-The ChartBlender system was motivated by our work experience on a data video design in collaboration with a prominent Chinese media organization.
-In this project, we were responsible for designing visualization charts and embedding the charts into a video.
-To support this task, we surveyed existing tools and systems for creating chart-embedded video.
-One notable system was VisTellAR [9], which enables end-users to record short videos and embed data visualizations using augmented reality (AR) techniques.
-VisTellAR tracks static elements in the video and anchors charts accordingly, offering a lightweight and mobile-friendly approach.
-Although VisTellAR is not open-source, we introduced it to our collaborators as a representative example of chart-embedded video tools.
-They noted, however, that end-users seldom embed charts into videos directly.
-In contrast, professional data journalists (our target users) require such capabilities but demand far greater precision and control.
-They highlighted three critical requirements: (1) accurate video motion tracking that can handle both static and dynamic objects; (2) compatibility with diverse video sources, including those lacking camera metadata, as is common in real-world media production; and (3) support for multiple dynamic charts within a single video, with fine-grained control over timing and placement to maintain contextual relevance and visual coherence.
-We observed a significant gap in tools that enable seamless and efficient embedding of visualizations into real-world videos.
-In current practice, designers often resort to manually inserting pre-rendered charts frame by frame using software like Adobe After Effects.
-This process so labor-intensive that a ten-second clip can require an entire day of work according to our own experience (see the example video  [^1] http://edu.people.com.cn/n1/2023/0915/c1006-40078200.html (2’44”, 3’25”) that we produced in this way).
-The majority of this effort goes into repetitive adjustments of position, scale, and orientation, rather than creative or analytical tasks.
-This inefficiency underscores the need for a dedicated authoring tool that can automate chart-video synchronization.
-Through close discussions with our collaborators, we distilled a set of core design requirements that guided the development of ChartBlender.
+In this section, we present the design of ChartBlender.
+We first introduce our collaboration with a national Chinese media organization.
+Drawing on their feedback and our observations, we distill four design requirements (R1–R4) and then provide an overview of the system.
+3.1 Collaboration with Professional Data Journalists
+The development of ChartBlender was motivated by a collaboration with a national Chinese media organization.
+Our research team was invited to provide technical support for a micro-documentary on the College Student Volunteer Service Western Project, which aimed to narrate volunteers’ stories and the program’s long-term impact on regional development.
+We assembled a project team of two professors and four graduate students with expertise in data visualization.
+We worked closely with five professional data journalists from the organization (average 15 years of work experience), who were highly familiar with data videos.
+Using raw footage and data provided by the organization, we were tasked with producing three chart-embedded video segments that required high-precision visual integration.
+One segment embedded a 3D bar chart into an aerial drone shot to depict year-by-year changes in volunteer participation, with the chart appearing anchored to the terrain.
+The other two segments overlaid key indicators that had to remain tightly aligned with moving targets in the footage.
+The collaboration followed a 15-day production cycle with daily online review meetings and rapid iterative revisions.
+In each meeting, our team presented the latest composites together with intermediate project files, and the journalists evaluated them shot by shot.
+They requested targeted adjustments such as changing anchoring targets, fine-tuning entrance and exit timing, refining visual styling for on-screen legibility, and repositioning overlays.
+Between meetings, we implemented these requests and circulated updated renders for asynchronous feedback, which was then consolidated and addressed in subsequent reviews.
+This recurring loop of authoring, critique, and revision not only ensured visual precision but also exposed where existing workflows break down, what corrections experts repeatedly perform, and which controls they consider essential for professional-grade results.
+To support the task, we surveyed prior systems for chart-embedded video authoring.
+One notable system is VisTellAR [VisTellAR], which enables end users to capture short videos and embed data visualizations through augmented reality (AR).
+VisTellAR tracks static elements in the scene and anchors charts accordingly, offering a lightweight, mobile-friendly workflow.
+Although VisTellAR is not open source, we introduced it to our collaborators as a concrete reference for discussing chart-in-video authoring.
+They noted, however, that unlike end-user workflows, professional data journalists (our target users) typically embed charts during post-production and demand substantially higher precision and finer-grained control.
+Across production iterations and review discussions, three needs repeatedly surfaced: (1) accurate motion tracking that supports both static structures and dynamic objects; (2) compatibility with diverse footage, including videos without camera metadata, which is common in practice; and (3) support for multiple dynamic charts within a single video, with fine-grained control over timing and placement to maintain visual coherence.
+In the absence of dedicated tools, we followed a typical professional workflow to produce the segments.
+We first created charts in 3D software, exported them as 3D assets, imported the assets into Adobe After Effects, applied automatic motion tracking to obtain an initial alignment, and then refined the results by adjusting position and orientation frame by frame before final compositing.
+During this process, we observed several bottlenecks that hinder seamless and efficient embedding.
+Although professional compositing software provides tracking functions, tracking results can be unstable in challenging shots, requiring repeated adjustments of position, scale, and orientation to maintain spatiotemporal consistency.
+Moreover, chart authoring and motion tracking are decoupled, preventing designers from previewing and refining charts in context while editing.
+In our experience, producing a 10-second clip with this workflow can take several hours, even a full workday, with much of the effort devoted to repetitive refinements rather than higher-level design work.
+These observations motivate a dedicated authoring tool that tightly integrates chart design with robust tracking to automate chart–video synchronization while preserving the precise control required in professional production.
+3.2 Design Requirements
+Through discussions with our collaborators, we distilled four design requirements (R1-R4) that guided the development of ChartBlender, which shaped both the system’s architecture and interface.
   • R1
@@
 Visualizations should be embedded into the video by aligning with camera depth and adopting a perspective consistent with the video scene.
-Such alignment ensures that the visualizations appear perceptually native to the scene, rather than as artificial post-production overlays.
+Such alignment ensures that the visualizations appear seamless to the scene, rather than as artificial post-production overlays.
   • R3
@@
 As the video plays, the visualizations should move in synchrony with the video, including the movement of both the camera and objects in the scene.
-To achieve this synchronization, it is necessary to track camera and object motion accurately.
+To achieve this synchronization, it is necessary to accurately track both camera and object motion.
   • R4
@@
 4 Chart Embedder
-In this section, we present a tracking algorithm for synchronizing visualizations with video motion.
-The algorithm supports two scenarios: (1) keeping visualizations stably anchored to the background as the camera moves (camera tracking), and (2) allowing visualizations to follow moving objects within the scene (object tracking).
+The Chart Embedder module supports two tracking pipelines: camera tracking and object tracking.
+Camera tracking keeps an embedded 3D visualization spatially consistent with a static scene element throughout the video (Fig.
+3(a)), whereas object tracking keeps the visualization aligned with a moving target (Fig.
+3(b)).
 As shown in Fig.
-2, given an input video consisting of $N$ frames and a user-specified target point $(x,y)$, our goal is to generate a sequence of 3D pose matrices $O_{t}$.
-Each pose matrix represents the position and rotation of the tracked point in 3D space at frame $t$.
-It begins with depth estimation (Fig.
-2 (a)) from the input video frames, which produces per-frame depth maps.
-These maps are used in two parallel modules: camera tracking (Fig.
-2 (b)), which computes frame-to-frame transformations and accumulates them into a global camera pose; and object tracking (Fig.
-2 (c)), which tracks a user-specified point across frames and combines its 2D trajectory with depth information to estimate a smoothed 3D object pose.
-Finally, the estimated object pose is applied to the visualization, which is then overlaid onto the video scene.
-Figure 2: The pipeline of the Chart Embedder.
-4.1 Depth Estimation
-Our system employs a monocular depth estimation model, Metric3D[38], to generate per-frame metric depth maps directly from original video frames.
-The major advantages of adopting Metric3D are twofold: (1) it predicts absolute metric depth without requiring ground-truth depth or camera intrinsics, enabling zero-shot deployment on in-the-wild videos; (2) it achieves high generalization capability across scenes, owing to its large-scale training on diverse data.
-Given these benefits, we apply Metric3D to each of the $N$ video frames to generate a sequence of dense depth maps $D_{n}$, $n\in{0,1...,N-1}$.
-For a user-specified 2D point $(x,y)$, we can extract its first frame’s depth value from the depth map as $d=D_{0}(x,y)$.
-This depth information serves as a crucial foundation for the subsequent camera tracking and object tracking processes, enabling accurate spatial reasoning in both modules.
+4, both pipelines start with (A) depth estimation, which provides per-frame depth cues needed to lift 2D image measurements into 3D and to render charts with perspective-consistent projection.
+They then differ in their motion estimation stage: (B1) camera pose estimation for camera tracking recovers the camera trajectory so a chart can remain anchored to a fixed location in the scene despite camera motion, whereas (B2) object trajectory estimation for object tracking recovers the 3D motion of a target so a chart can follow the object over time.
 Figure 3: An illustration of the two motion embedding modes supported by our system: Camera Tracking and Object Tracking.
@@
 In contrast, Object Tracking binds the visualization to a moving object, attaching a chart to a galloping horse.
-4.2 Camera Tracking
-In ChartBlender, camera tracking ensures that embedded 3D visualizations remain spatially consistent with the static element throughout the video (Fig.
-3(a)).
-To achieve this, we estimate the camera’s motion through a three-step process.
-First, we perform (1) relative pose estimation to compute the geometric transformation between adjacent video frames, capturing how the camera moves over time.
-This information is then used in (2) camera pose estimation, accumulating relative transformations to recover the camera’s trajectory in the scene.
-Finally, we conduct (3) 3D pose recovery to infer the camera’s position and orientation in 3D space, enabling visualizations to be accurately projected.
+Figure 4: The two pipelines of the Chart Embedder.
+The camera tracking pipeline consists of (A) depth estimation followed by (B1) camera pose estimation.
+The object tracking pipeline consists of (A) depth estimation followed by (B2) object trajectory estimation.
+4.1 Depth Estimation
+Unlike general-purpose 2D trackers that return image-plane trajectories, chart embedding requires 3D, scene-referenced motion signals: we need depth to lift user-selected points from pixels into 3D space and to compute poses that keep a chart perspective-consistent and spatially stable under camera motion.
+To obtain the required spatial context, our system uses a monocular depth estimation model, Metric3D [yin2023metric], to predict metric depth maps for each video frame.
+We adopt Metric3D for two reasons: (1) it predicts absolute metric depth without requiring ground-truth depth or camera intrinsics, enabling zero-shot deployment on in-the-wild videos; and (2) it generalizes well across diverse scenes due to large-scale training data.
+Given these benefits, we apply Metric3D to each of the $N$ frames to obtain a sequence of dense depth maps ${D_{n}}_{n=0}^{N-1}$.
+For a user-specified 2D point $(x,y)$, we extract the depth value in the first frame as $d=D_{0}(x,y)$.
+This depth signal provides the foundation for subsequent camera and object tracking, enabling spatial reasoning in both modules.
+4.2 Camera Pose Estimation
+In ChartBlender, camera pose estimation aims to recover a consistent camera trajectory across frames so that a visualization anchored to a static scene location remains stable under camera motion.
+We decompose this problem into three steps.
+First, we perform (1) relative pose estimation to compute the geometric transformation between adjacent frames, capturing local camera motion that can be robustly inferred from short temporal baselines.
+Second, we perform (2) global pose estimation by accumulating these relative transformations to obtain a per-frame camera pose in a shared coordinate system, which is required for world-consistent anchoring over time.
+Finally, we perform (3) 3D pose recovery to lift the user-specified 2D anchor (with predicted depth) into a world-space reference and derive the corresponding per-frame chart pose, enabling accurate 3D placement and projection throughout the video.
 Relative pose estimation.
@@
 These pairwise transformations form the foundation for reconstructing the full camera trajectory.
-We employ an RGBD odometry model provided by Open3D [39, 40] to compute the relative camera motion between adjacent RGBD frames.
+We employ an RGBD odometry model provided by Open3D [Zhou2018, park2017colored] to compute the relative camera motion between adjacent RGBD frames.
 Given two consecutive frames $n$ and $n+1$, RGBD odometry estimates a transformation matrix.
 Specifically, we extract a transformation matrix $T_{n\rightarrow n+1}$ that aligns the RGB and depth images.
-Each transformation is a $4\times 4$ homogeneous transformation matrix composed of a rotation matrix $R_{n\rightarrow n+1}$ and a translation vector $t_{n\rightarrow n+1}$:
-$T_{n\rightarrow n+1}=\left[\begin{array}[]{cc}R_{n\rightarrow n+1}&t_{n\rightarrow n+1}\\ 0&1\end{array}\right]$ (1)
-Camera pose estimation.
+Each transformation composed of a rotation matrix $R_{n\rightarrow n+1}$ and a translation vector $t_{n\rightarrow n+1}$.
+Global pose estimation.
 While relative poses describe local motion between frames, embedding visualizations in a stable, world-aligned coordinate system requires each frame’s absolute camera pose.
-This enables consistent rendering of the 3D object across the video sequence.
-We compute the absolute camera pose $M_{n}$ at each frame $n$ by chaining all previous relative transformations starting from the first frame:
-$M_{n}=T_{1\rightarrow 2}T_{2\rightarrow 3}\cdots T_{{n-1}\rightarrow n}$ (2)
-This accumulation step reconstructs the global camera trajectory, allowing us to express each frame’s viewpoint in the same coordinate system.
+This enables a consistent rendering of the 3D object throughout the video sequence.
+We compute the absolute camera pose $M_{n}$ at each frame $n$ by chaining all previous relative transformations starting from the first frame: $M_{n}=T_{1\rightarrow 2}T_{2\rightarrow 3}\cdots T_{{n-1}\rightarrow n}$ This accumulation step reconstructs the global camera trajectory, allowing us to express each frame’s viewpoint in the same coordinate system.
 3D pose recovery.
-To render the 3D visualization as if it were physically anchored in the scene, we must compute its position in the global world coordinate system.
-Given a user-specified 2D point $(x,y)$ in the image and its corresponding predicted depth $d$, we first back-project the point into the 3D camera coordinate system using the inverse of the camera intrinsic matrix $K^{-1}$.
-The resulting 3D point in the camera frame is:
-$P_{\text{n}}=K^{-1}\cdot\left(d_{0}\cdot\tilde{o}_{n}\right)$ (3)
-where $P_{n}$ denotes the 3D position of a tracked point at frame $n$, $\tilde{o}_{n}=[x,y,1]^{T}$ is the homogeneous pixel coordinate.
+To render the 3D visualization as if it were physically anchored in the scene, we should compute its position in the global world coordinate system.
+Given a user-specified 2D point $(x,y)$ in the image and its predicted corresponding depth $d$ predicted by the depth estimation module, we first reproject the point back into the 3D camera coordinate system using the inverse of the intrinsic matrix of the camera $K^{-1}$.
 Next, we transform this point into the world coordinate system using the inverse of the absolute camera pose $M_{n}$, computed from accumulated relative poses.
-The object’s global pose $O_{n}$ is then given by:
-$O_{n}=M_{n}^{-1}\cdot P_{\text{n}}$ (4)
 This transformation places the 3D visualization at a consistent location in the world coordinate system, regardless of camera motion.
 It ensures that the visualization remains spatially stable and visually aligned with the static background in the video.
-4.3 Object Tracking
-In ChartBlender, the goal of object tracking is to enable visualizations to remain spatially aligned with moving elements in the video scene (Fig.
-3(b)).
-To achieve this, we estimate the 3D motion of target objects from monocular video input through a three-stage pipeline.
-First, we perform (1) 2D point tracking to extract the object’s pixel-wise trajectory across frames, capturing how it moves within the image plane.
-This trajectory serves as the basis for the second stage, (2) 3D translation estimation, which lifts the 2D motion into 3D space to approximate the object’s position relative to the camera over time, allowing visualizations to maintain consistent placement as the object moves.
-Finally, we perform (3) 3D rotation estimation to recover the object’s orientation changes, ensuring that embedded visualizations can rotate coherently with the object’s local motion, further enhancing the perceptual integration.
+4.3 Object Trajectory Estimation
+Object trajectory estimation aims to recover a per-frame 3D pose for a moving target so that an attached visualization follows the object with stable position and coherent orientation over time.
+We decompose this problem into three stages.
+Specifically, we estimate per-frame object poses via (1) 2D point tracking, which extracts a pixel trajectory; (2) 3D translation estimation, which combines the tracked 2D positions with the predicted depth to recover the object’s 3D displacement; and (3) 3D rotation estimation, which infers a smooth orientation change to keep the visualization consistently aligned with the object’s motion.
+Together, these stages produce per-frame poses that drive motion-synchronized chart embedding for dynamic targets.
 2D point tracking.
-We begin by identifying salient points on dynamic objects and tracking their 2D positions across frames.
-To accomplish this, we use BootsTAP [41], a high-precision keypoint tracking algorithm that leverages dense visual correspondence estimation.
-For each target point $(x,y)$, the tracker generates a pixel-level trajectory over time, denoted as $p_{n}=(u_{n},v_{n})\rightarrow p_{n+1}=(u_{n+1},v_{n+1})$.
-The 2D trajectories capture the apparent motion of object features in image space and form the foundation for estimating real-world displacement in 3D, which is critical for subsequent motion reconstruction.
+We identify salient points on the target object and track their 2D locations across frames.
+We use BootsTAP [doersch2024bootstap], a high-precision keypoint tracker based on dense visual correspondence.
+For a target point $(x,y)$, the tracker outputs a pixel-level trajectory $p_{n}=(u_{n},v_{n})$ across frames.
+These trajectories capture apparent motion in image space and serve as the basis for recovering 3D displacement.
 3D translation estimation.
-To reconstruct the spatial motion of tracked points, we lift the 2D trajectories to 3D space.
-For each frame, the 2D trajectory is projected into 3D space using the depth map of the tracked point in the camera coordinate system.
-$d_{n}=D_{n}(p_{n})$ (5)
-However, monocular depth estimates are inherently noisy and temporally unstable due to factors such as low-texture regions, occlusions, and variations in lighting.
-These inconsistencies can lead to jitter and discontinuities in the reconstructed 3D trajectories, especially when tracking fine-grained object motion.
-Such artifacts degrade both the physical plausibility and visual coherence of the motion path, undermining downstream tasks like 3D alignment and visualization anchoring.
-To address this, we apply temporal smoothing to the depth signal using quadratic regression over a sliding window:
-$\hat{d_{n}}=\text{QuadraticRegression}(d_{n})$ (6)
-As previously defined in Equation (3), we recover the 3D translation vector $t_{n}$ of the target point in the camera coordinate system by back-projecting its 2D image coordinate using the estimated depth and the intrinsic matrix K.
+We lift 2D trajectories into 3D by combining them with per-frame depth.
+For each frame, we obtain the depth at the tracked pixel from the corresponding depth map, $d_{n}=D_{n}(p_{n})$, and back-project $p_{n}$ to a 3D point $P_{n}$ in camera coordinates.
+Because monocular depth can be noisy and temporally inconsistent (e.g., due to low-texture regions, occlusion, or lighting changes), the recovered 3D trajectory may exhibit jitter.
+We therefore smooth the depth signal over time using quadratic regression within a sliding window, and then compute the per-frame translation vector $t_{n}$ via back-projection with the smoothed depth and intrinsic matrix $K$.
 3D rotation estimation.
-Finally, we estimate object rotation based on the direction of motion in 3D space to ensure that visualizations remain properly oriented when attached to moving objects.
-Specifically, we define the per-frame motion vector as $V_{n}=P_{n+1}-P_{n}$.
-We obtain the rotation vector $R_{n}$ by constraining the object’s local X axis to follow $V_{n}$.
-To support stable rotation estimation, we first regularize the raw 3D point trajectories by minimizing sudden changes in displacement.
-This is implemented as a second-order temporal smoothing loss:
+We estimate object rotation from the direction of motion in 3D so that the visualization remains properly oriented as the object moves.
+We define the per-frame motion vector as $V_{n}=P_{n+1}-P_{n}$, and obtain a rotation $R_{n}$ by constraining the object’s local $x$-axis to align with $V_{n}$.
+To stabilize rotation, we first regularize 3D point trajectories by minimizing sudden changes in displacement using a second-order temporal smoothing objective:
 $\min_{\hat{P}}\sum_{n}\|\hat{P}_{n+1}-\hat{P}_{n}\|^{2}+\lambda\sum_{n}\|\hat{P}_{n+2}-2\hat{P}_{n+1}+\hat{P}_{n}\|^{2}$ (1)
-Next, we apply the same principle in the velocity domain to ensure that the direction of motion also varies smoothly over time. The corresponding velocity regularization is formulated as:
-$\mathcal{L}(V_{n})=\sum_{n}\|V_{n+1}-V_{n}\|^{2}+\lambda\sum_{t}\|V_{n+1}-2V_{n}+V_{n-1}\|^{2}$ (8)
-This second-order smoothing constraint encourages both consistent motion direction and smooth acceleration profiles, mitigating jitter and improving temporal coherence in the estimated trajectories.
-Given the smoothed translation vector and rotation vector, the final object pose at each frame is defined as $O_{n}=\{R_{n}\mid t_{n}\}$.
+where $\lambda$ controls the strength of the acceleration penalty.
+We apply the same principle in the velocity domain to encourage smooth motion direction over time, mitigating jitter and improving temporal coherence.
+Given the estimated translation $t_{n}$ and rotation $R_{n}$, the per-frame object pose is defined as $O_{n}=\{R_{n}\mid t_{n}\}$.
 5 ChartBlender
@@
 The user interface, as shown in Fig.
-4, consists of three major views: the Chart View (Fig.
-4-1), the Video View (Fig.
-4-2), and the Timeline View (Fig.
-4-3).
+5, consists of three major views: the Chart View (Fig.
+5-1), the Video View (Fig.
+5-2), and the Timeline View (Fig.
+5-3).
 In the Chart View, users can upload a dataset, after which the raw data is displayed directly within the view.
@@
 5.2 Visualization Design
-Our goal is to design visualizations that are well-suited for integration within videos.
-This requires rethinking how charts are structured and styled to maintain both visual clarity and contextual harmony in videos[42].
-To this end, we collected a total of 200 high-quality videos and manually segmented them into 502 clips, each containing at least one embedded visualization.
-As shown in Fig.
-6, these videos span ten diverse domains.
-We also analyzed the most commonly used chart types across the dataset.
+We observed that visualizations appearing in videos often differ from conventional charts.
+To understand how embedded visualizations are designed in practice and to inform the design of our chart library, we followed a systematic workflow: we collected a real-world video corpus and synthesized recurring strategies into a design space of embedded-visualization principles, offering practical implications for both researchers and designers.
+5.2.1 Methods
+To ground our design in practice, we assembled a corpus of 200 publicly available, high-quality videos featuring embedded visualizations from online platforms (e.g., YouTube and Bilibili) and official outlets in domains such as news, sports, technology, and education.
+Each video was manually segmented at visualization entry and exit points, yielding 502 clips containing at least one embedded visualization.
+Two trained annotators coded each clip using a codebook that combined deductive categories (chart type, 2D/3D representation, spatial relation to the scene, motion coupling to camera/object, occlusion handling, and labeling/styling) with inductive categories for emergent patterns.
+After an initial calibration phase, disagreements were resolved through discussion.
+We also recorded domain metadata spanning ten categories (e.g., economy, sports, ecology, and transportation), as summarized in Fig.
+7.
+Figure 6: Design space for visualizations suitable for embedding in video, which illustrates five representative visualization forms combined with two motion anchoring strategies.
+It reveals common design patterns for how visualizations are structured and integrated within dynamic video scenes across diverse real-world applications.
+Figure 7: Distribution of chart types and video domains.
 Furthermore, we identified several recurring design patterns in how visualizations are embedded within real-world video content.
 These patterns are summarized as follows.
-Figure 5: Design space for visualizations suitable for embedding in video, which illustrates five representative visualization forms combined with two motion anchoring strategies (object tracking and camera tracking).
-It reveals common design patterns for how visualizations are structured and integrated within dynamic video scenes across diverse real-world applications.
-Figure 6: Distribution of chart types and video domains.
-First, a majority of charts (130 cases) incorporate background to improve readability, with a subset (21) employing semi-transparent backgrounds to maintain visual access to the underlying scene.
-Notably, gridlines were selectively applied to support data interpretation, with horizontal guides appearing more frequently (109) than vertical ones (43), while 64 charts omitted them entirely.
-Second, very few charts (only 4) included legends, indicating a general preference for minimalist presentations in time-constrained video contexts.
-Nevertheless, essential structural components such as axes (retained in 171 charts) and titles (present in 168 charts) were commonly preserved.
-Third, when visualizations were associated with specific objects in the scene, designers frequently employed visual linking strategies to establish a clear connection between the chart and its target (68).
-These strategies included directional cues such as arrows, lines, and bounding boxes.
-The choice of pointer type often varied depending on scene complexity: in simple scenes with few moving objects, minimal indicators such as triangles were typically sufficient; in more complex scenes with multiple overlapping motions, designers commonly used bounding boxes or anchored pointers to more clearly distinguish the tracked target.
+(1) Legibility.
+Background panels are widely adopted to enhance readability (130/502), with a minority employing semi-transparent panels to maintain visual context (21/502).
+Gridlines are used selectively: horizontal guides appear more frequently (109) than vertical ones (43), while a notable number of charts omit gridlines entirely (64).
+(2) Referential Linking.
+When visualizations correspond to specific objects in the scene, designers often employ visual linking techniques (e.g., arrows, lines, or bounding boxes) to establish clear referential ties (68).
+The choice of pointer type tends to adapt to scene complexity: in simple scenes with sparse motion, minimal markers like triangles often suffice; in cluttered environments with overlapping movements, bounding boxes or anchored pointers are preferred to unambiguously identify the target.
+(3) Simplicity.
+Legends are rarely used (only 4 instances), reflecting a preference for minimalism in video formats.
+In contrast, essential structural elements such as axes (retained in 171 charts) and titles (present in 168 charts) are commonly preserved to maintain baseline interpretability.
+5.2.2 Design Space
 These findings highlight a set of implicit yet widely adopted design patterns for video-suited visualizations design, which we synthesize into a structured design space, as illustrated in Fig.
-5.
-This design space is organized along two key dimensions: the motion mode and the visualization type.
-It distinguishes between two modes of motion (object tracking and camera tracking) corresponding to whether the visualization follows a moving target in the scene or remains fixed relative to the global environment.
-Horizontally, it categorizes common visualization types, including text annotations, line charts, area charts, bar charts, and 3D charts, as well as their corresponding linking and layout strategies (e.g., aiming lines, bounding boxes, or indicators).
-Based on this design space, we developed a visualization chart library using Three.js, providing a collection of chart templates optimized for embedding in dynamic video scenes.
+6.
+Notably, although we identified a broad spectrum of ten chart types across diverse application scenarios (as shown in Fig.
+7), we focused the design space on the five most commonly used chart types in the corpus to maximize generalizability.
+The space is organized along three primary dimensions: (i) chart type, covering five common forms, including data, text, line charts, bar charts, and area charts; (ii) motion mode, including object-tracking and camera-tracking embeddings; and (iii) temporal behavior shown at the bottom of each schematic.
+We describe each visual element by its on-screen duration and animation scheme.
+Importantly, these dimensions are not independent; instead, they co-vary with production constraints and communicative intent in systematic ways.
+First, the choice of motion mode usually reflects what the visualization refers to.
+When the message is about a specific entity in the scene (e.g., a person, a vehicle, or a building), creators tend to use object-tracking so the overlay stays visually attached to the target.
+In these cases, the visualization is often kept compact (short text/data, small charts) and paired with explicit cues such as arrows or boxes.
+In contrast, when the visualization summarizes the scene or provides background context (e.g., overall trends or comparisons), creators more often use camera-tracking overlays.
+This allows larger chart panels and reduces the risk that the overlay jitters or drifts with local object motion.
+Second, temporal behavior is largely driven by attention and editing rhythm.
+Embedded charts rarely stay static on screen.
+Instead, they appear when the narration or shot content calls for data, remain long enough to be read, and then disappear to return attention to the footage.
+As a result, the most common animations are subtle and functional (e.g., fade-in/out or simple unveil), because aggressive animation competes with motion in the video and makes reading harder.
+Finally, these observations explain why a design space is useful for system building.
+In professional workflows, creators repeatedly adjust the same small set of things: what the overlay is anchored to, how strictly it tracks, where it sits to avoid covering important content, how strong the background panel should be for legibility, and when it enters/exits.
+The design space guided our system’s chart library design by turning the most frequently adjusted parameters into explicit controls and offering presets for common embedded chart styles.
 5.3 Interaction
-To support efficient visualization authoring and seamless synchronization within video, we propose a set of interactions in ChartBlender:
-Upload. Users can upload datasets and video files either by clicking the upload button or by dragging and dropping files into the designated panel.
-Data Mapping. Data attributes can be mapped to visual channels through direct manipulation: users drag data fields onto chart elements to define encoding relationships.
-Zoom and Pan. Users can interactively adjust the visualization view: left-click and drag to reposition charts, scroll to zoom in and out, and right-click and drag to rotate.
-Chart and Style Switching.
-Users can switch between different chart types via tabs in the interface.
-They can also adjust visual styling, such as color schemes, by clicking style options or using a color picker tool.
-Anchor Selection.
-To ensure accurate chart placement, users can click on static or dynamic objects in the video to define visual anchor points.
-These anchors are used to bind visualizations to specific elements in the scene, supporting consistent spatial alignment across frames.
-Track Editing.
-Users can manage chart tracks via the timeline view.
-Tracks can be added or removed by clicking the corresponding buttons in the toolbar on the left.
-On the right, users can adjust the timing and duration of each chart by dragging chart tracks directly along the timeline.
+To support efficient authoring and reliable synchronization, ChartBlender provides an integrated workflow that covers data binding, spatial anchoring, and temporal configuration.
+Key interaction techniques include:
+Upload and Data Mapping.
+Users upload video footage and datasets via drag-and-drop or file selection.
+When timestamps are available, the system aligns data records with the video timeline.
+Users map data fields to visual channels by dragging them onto chart components, and the preview updates immediately to reflect mapping changes.
+One-click Tracking Anchor.
+Users specify a tracking anchor by simply clicking a pixel in the video frame.
+Based on the selected motion mode, the system either (1) attaches the visualization to a moving target via Object Tracking or (2) keeps the chart anchored to the static scene via Camera Tracking in world coordinates.
+3D Adjustment.
+Users pan, zoom, and rotate the chart in 3D to fine-tune placement: dragging pans, scrolling zooms, and right-dragging rotates the view.
+These operations work together with anchoring to achieve stable alignment and perspective-consistent embedding.
+Real-time Preview.
+ChartBlender provides continuous, in-context preview throughout authoring.
+While editing, users can immediately see changes to chart styling and 3D viewpoint directly on the video.
+After rendering, users can play back the result to preview the chart’s motion in real time and verify alignment and visual coherence over time.
+Multi-chart Editing.
+Users manage when each visualization appears using the timeline.
+Multiple charts are organized as separate tracks and can be added or deleted with simple timeline operations.
+Users adjust a chart segment’s entry, exit, and duration by dragging its endpoints, and selecting a segment brings that chart into focus for independent editing of its data mapping, style, and placement.
 6 Evaluation
-We evaluated the ChartBlender and the key techniques through three approaches: (1) a quantitative experiment to assess the precision of camera tracking, (2) a quantitative experiment to assess the precision of object tracking, and (3) domain expert interviews to evaluate the overall usability of the system.
-6.1 Camera Tracking Precision
-In our system, camera tracking is implemented using RGBD odometry, which estimates frame-by-frame camera motion by leveraging both RGB appearance and depth information.
-As a result, the quality of depth estimation plays a critical role in the precision of camera tracking.
-To evaluate this dependency, we compare how different monocular depth models affect the overall pose estimation performance of the RGBD odometry pipeline.
+We evaluated ChartBlender through (1) two quantitative experiments that assess the precision of our two tracking pipelines and (2) interviews with domain experts that assess system usability.
+6.1 Evaluation of Two Tracking Pipelines
+Our system relies on two tracking pipelines: camera tracking and object tracking.
+The camera tracking pipeline consists of (A) depth estimation followed by (B1) camera pose estimation.
+The object tracking pipeline consists of (A) depth estimation followed by (B2) object trajectory estimation.
+The two pipelines share the same depth estimation stage.
+However, to the best of our knowledge, there’s no open-source system that provides an end-to-end baseline for motion-synchronized chart embedding.
+We therefore construct baselines using a component substitution strategy: we replace individual stages with representative models identified in our literature review and evaluate the resulting combinations under identical experimental settings.
+This strategy allows us to quantify the tracking precision of two pipelines.
+6.1.1 Precision of the Camera Tracking Pipeline.
 Experimental Setup.
-To evaluate the impact of depth estimation quality on camera pose recovery, we conducted a comparative experiment using different monocular depth models combined with RGBD odometry.
-Specifically, we compared Metric3D, which emphasizes metric depth precision, with ZoeDepth, which focuses on generalization and temporal consistency.
-We refer to the two pipelines as Metric3D + RGBD odometry (our proposed method) and ZoeDepth + RGBD odometry (baseline).
-We selected a subset of the DriveTrack dataset [43], consisting of 18 video sequences (1,806 frames in total), where camera motion was sufficiently pronounced and affected by dynamic foreground objects.
-For each method, predicted depth maps were fed into RGBD odometry implemented in Open3D [39], which uses a hybrid geometric-photometric optimization [40] to estimate per-frame camera extrinsics.
-A fixed intrinsic matrix of the form [1000, 1000, width/2, height/2] was used for all sequences.
+The camera tracking pipeline comprises two core stages: depth estimation and camera pose estimation.
+In ChartBlender, the camera tracking pipeline is implemented as Metric3D + Open3D RGBD odometry [Zhou2018, park2017colored].
+To validate the effectiveness of our pipeline, we construct a baseline by substituting the depth estimation component, yielding ZoeDepth + RGBD odometry (baseline).
+We choose this baseline because Open3D RGB-D odometry is a frequently used and stable open-source camera pose estimator, making it a reliable and fixed module.
+ZoeDepth [bhat2023zoedepth] is another top-performing depth estimator identified in our literature review.
+The goal of our experiment is to compare whether our depth estimation model (Metric3D) achieves better performance than ZoeDepth when paired with the same tracking module.
+We evaluate on the subset of DriveTrack [koppula2024tapvid] (18 sequences, 1,806 frames) with ground-truth camera poses, where camera motion is sufficiently pronounced, and scenes include dynamic foreground objects.
+These conditions closely resemble the real-world videos in which our system embeds charts.
+For each frame, the predicted depth map is fed into RGBD odometry to estimate camera poses.
 Evaluation Metric.
-To quantitatively evaluate the precision of our camera tracking approach and assess the influence of different depth estimation models, we measured the geometric discrepancy between the estimated camera poses and the ground-truth extrinsics provided by the TAPVid-3D dataset [34].
-For each frame, we compared the predicted camera pose $(R_{\text{pred}},t_{\text{pred}})$ with the corresponding ground-truth pose $(R_{\text{gt}},t_{\text{gt}})$, and computed both rotation and translation errors.
-The rotation error was defined as the angular difference between the predicted and ground-truth orientations.
-Specifically, we first computed the relative rotation matrix $\Delta R=R_{\text{gt}}^{-1}R_{\text{pred}}$, and then calculated the rotation error as the angle of this rotation:
-$\theta=\arccos\left(\frac{\operatorname{trace}(\Delta R)-1}{2}\right),$ (9)
-This formulation captures the minimum angular difference between the two rotation matrices in 3D space.
-The translation error was defined as the Euclidean distance between the predicted and ground-truth camera positions:
-[…]
-Together, these two metrics quantify both orientation and positional discrepancies in the estimated camera trajectories.
-To evaluate overall precision and robustness to outliers, we report both the mean and median rotation and translation errors computed across all frames in the dataset.
+We report rotation and translation errors between predicted poses $(R_{\text{pred}},t_{\text{pred}})$ and ground truth $(R_{\text{gt}},t_{\text{gt}})$.
+Rotation error is computed from the relative rotation $\Delta R=R_{\text{gt}}^{-1}R_{\text{pred}}$ as the corresponding angular difference, and translation error is the Euclidean distance $\lVert t_{\text{gt}}-t_{\text{pred}}\rVert$.
 Results and Analysis.
-As shown in Table I, Metric3D consistently outperformed ZoeDepth in both rotation and translation precision when paired with RGBD odometry, achieving lower mean and median errors across all frames.
-Although both methods operate under the same fixed intrinsic setting, Metric3D provided more reliable metric depth, enabling more accurate frame-by-frame camera pose estimation.
-Another key reason for adopting Metric3D is its robustness in unconstrained environments where intrinsic calibration is unavailable, a common challenge in user-generated or in-the-wild videos.
-Despite the use of an approximate intrinsic matrix, the Metric3D + RGBD odometry pipeline maintained strong performance.
+Table I shows that our pipeline yields lower rotation and translation errors than the baseline (mean and median).
+Since both variants use the same odometry backend and camera intrinsics, the improvement is attributable to the depth backbone, indicating that Metric3D provides depth estimates that better support downstream pose recovery in our pipeline.
 TABLE I: Comparison on camera tracking. (Camera pose estimation errors)
@@
 ZoeDepth + RGBD odometry 10.544 11.061 7.598 7.610
-6.2 Object Tracking Precision
-We evaluated our system’s object tracking techniques by assessing the precision of the reconstructed 3D motion trajectories for dynamic objects.
-This reflects how well the system can estimate object movement in physical space.
+6.1.2 Precision of the Object Tracking Pipeline.
 Experimental Setup.
-In our system, object tracking is implemented by combining 2D keypoint tracking with monocular depth estimation to recover the 3D trajectories of dynamic objects across time.
-This process is essential for enabling visualization alignment with moving targets in the scene.
-The overall tracking precision depends on the quality of both the 2D keypoint trajectories and the estimated depth maps used for lifting to 3D.
-To evaluate this component, we conducted a systematic comparison of different tracking configurations by varying the underlying 2D tracking and depth estimation models.
-Specifically, we considered two keypoint tracking models, TAPIR [44] and BootsTAPIR [41], and two depth estimation models, ZoeDepth and Metric3D, resulting in four combinations in total.
-We performed this evaluation on a subset of the PStudio dataset [45], consisting of 36 video sequences (10,800 frames in total).
-These sequences feature clearly observable dynamic objects with prominent motion and minimal camera movement, allowing us to isolate object trajectory estimation from background motion.
-For each combination, 2D keypoints were first tracked across frames using the selected tracking model.
-These trajectories were then lifted into 3D space by back-projecting the 2D points with the corresponding depth estimates.
+Object tracking comprises two stages: depth estimation and object trajectory estimation.
+In ChartBlender, we implement this pipeline as Metric3D + BootsTAP, which is not the only reasonable choice.
+During our literature review, we also identified strong alternatives for each stage, including ZoeDepth[bhat2023zoedepth] for depth estimation and TAPIR[doersch2023tapir] for object trajectory estimation.
+To validate that our chosen pipeline remains the best-performing configuration under the same experimental conditions, we evaluate three additional baseline combinations: Metric3D + TAPIR, ZoeDepth + TAPIR, and ZoeDepth + BootsTAP.
+We performed this evaluation on a subset of the PStudio dataset [joo2015panoptic], which contains high-fidelity multiview recordings of social interaction scenes.
+We selected 36 video sequences (10,800 frames total) that include clearly visible dynamic objects undergoing complex 3D motion, while the camera remains mostly stationary.
+This setup allows us to isolate the quality of object trajectory reconstruction from the effects of camera movement.
 Evaluation Metric.
-We adopted the Average 3D Point Distance (APD3D) metric, introduced in TAPVid-3D [34], to evaluate the precision of 3D point tracking.
-This metric quantifies the proportion of predicted 3D points that lie within a specified Euclidean distance threshold of their corresponding ground-truth positions, averaged over time and across tracked points.
-A higher value reflects better spatial precision and temporal consistency of the reconstructed trajectories.
-Formally, the metric is defined as:
-$\mathrm{APD}_{3D}\equiv\frac{1}{V}\sum_{i,t}v_{t}^{i}\cdot\mathbf{1}\left(\lVert\hat{P}_{t}^{i}-P_{t}^{i}\rVert<\delta_{3D}(P_{t}^{i})\right),$ (11)
-where $\hat{P}_{t}^{i}$ and $P_{t}^{i}$ denote the predicted and ground-truth 3D positions of the $i$-th point at time $t$, $v_{t}^{i}$ is a visibility indicator, and $V$ is the total number of visible points.
-To quantify performance across different levels of precision, we report APD3D under a fixed set of distance thresholds: $[0.04,0.08,0.16,0.32,0.64]$.
+We adopted the Average 3D Point Distance (APD3D) metric [koppula2024tapvid], to evaluate the precision of 3D point tracking.
+APD3D measures the fraction of predicted 3D points that fall within a given Euclidean distance threshold of the corresponding ground-truth points, averaged over time and across tracks.
+Higher values indicate more accurate and temporally consistent 3D trajectories.
 Results and Analysis.
-As shown in TABLE II, BootsTAPIR consistently outperformed TAPIR across all configurations, confirming its superior capability in 2D temporal keypoint tracking.
-With regard to depth estimation, our primary objective is not to reconstruct metrically accurate 3D geometry, but to generate perspective-consistent visualizations that preserve spatial coherence between foreground and background elements.
-In this context, moderate depth errors are acceptable, provided that the overall depth structure remains stable and visually plausible.
+As shown in Table II, pipelines using BootsTAP consistently outperform those using TAPIR under the same depth backbone, indicating that BootsTAP provides more reliable 2D keypoint trajectories for downstream 3D reconstruction.
+Although ZoeDepth + BootsTAP achieves the highest APD3D score, we adopt Metric3D + BootsTAP in ChartBlenderfor better alignment with our camera tracking pipeline.
+Because both variants use the same 2D tracker (BootsTAP), the APD3D gap primarily reflects the choice of depth backbone.
+With regard to depth estimation, our objective is not metrically exact 3D trajectory recovery, but perspective-consistent chart embeddings that preserve spatial coherence between foreground and background elements.
+Therefore, moderate depth errors are acceptable, provided that the overall depth structure and recovered trajectories remain stable.
 […]
 […]
-BootsTAPIR + Metric3D (ours) 0.3424
-TAPIR + Metric3D 0.3416
-BootsTAPIR + ZoeDepth 0.3609
-TAPIR + ZoeDepth 0.3595
+Metric3D + BootsTAP (ours) 0.3424
+Metric3D + TAPIR 0.3416
+ZoeDepth + BootsTAP 0.3609
+ZoeDepth + TAPIR 0.3595
 Figure 8: Five example videos with embedded visualizations generated using ChartBlender
 6.2 Expert Interview
-To further evaluate the usability of ChartBlender, a series of interviews with five data journalists (denoted as E1-E5) were performed.
-They had over 5 years’ video editing experiences on average, and were very familiar with chart-embedded videos.
-All participants provided informed consent prior to the study.
+Participants.
+To further evaluate the usability of ChartBlender, we conducted semi-structured interviews with five data journalists (E1–E5; Male=2, Female=3; Age: 32-45).
+Two were from our collaborating media organization, while the remaining three were recruited from other newsrooms to broaden perspectives and reduce organizational bias.
+Participants had, on average, over ten years of video-editing experience and were familiar with chart-embedded videos.
+All participants provided informed consent before the study.
 Procedure.
 The interviews were performed via an online meeting system.
-Each interview started by a 10-minute introduction about the system.
-Participants were first introduced to our system through a brief tutorial delivered by authors, and were given two task:
+Each interview started with a 10-minute introduction about the system.
+Participants were first introduced to our system through a brief tutorial delivered by the authors and were given two tasks:
 Task 1 - Goal-oriented (10 minutes).
@@
 The result should consist of a bar chart that remains fixed in the field and an area chart that moves along with the vehicle.
-Task 2 - Open-ended (30 minutes). Participants are asked to create two novel data videos by choosing among 10 short video clips (10” duration on average).
-Finally, the experts were asked to finish a questionnaire, followed by an interview after they explored the functionalities of ChartBlender.
-The whole process lasted for about one hour, and the interviews were recorded for analysis.
+Task 2 - Open-ended (30 minutes). Participants are asked to create two novel data videos by choosing among 10 short video clips (10" duration on average).
+Measures.
+After the tasks, experts participated in a semi-structured interview.
+Each session lasted approximately 60 minutes.
+We gathered two data sources: (i) The completed videos, (ii) audio recordings of think-aloud protocols and interviews.
+Transcripts and observation notes were analyzed using a reflexive thematic analysis.
+Three authors independently performed open coding to generate candidate codes, reconciled discrepancies through discussion, and iteratively refined a shared codebook until reaching consensus on higher-level themes.
 Result. Fig. 8 presents sample videos generated by expert users using our system. In follow-up interviews, the experts offered a range of insightful comments, which are summarized as follows:
@@
 All experts agreed that our system was highly effective in supporting both camera tracking and object tracking.
-In particular, E1 commented that “This system can accomplish the work of several hours, or even days, in just two or three minutes.” E2 mentioned that “This system allows me to see the real-time effect of chart edits in the video… I no longer need to repeatedly switch between project files, which used to waste a lot of time.” E5 also noted that “The speed of result generation is so fast that I can quickly create multiple demos to observe where the chart embedding works best in the video.”
+In particular, E1 commented that “This system can accomplish the work of several hours, or even days, in just two or three minutes.” E2 mentioned that “This system allows me to see the real-time effect of chart edits in the video… I no longer need to repeatedly switch between project files, which used to waste a lot of time." E5 also noted that “The speed of result generation is so fast that I can quickly create multiple demos to observe where the chart embedding works best in the video."
 Accurate Results.
@@
 Specifically, they appreciated how seamlessly the system integrates chart creation and embedding into videos, accomplishing highly professional tasks (E1, 2, 5).
-E1 mentioned that “This system can complete highly professional tasks without requiring extensive skills or knowledge.” E5 noted that “the chart can be translated and rotated in 3D, and it integrates seamlessly with the video.” They were also satisfied with the chart creation aspect.
-E4 noted that “the data mapping feature eliminated the need for selecting data and creating visualizations in 3D software.” E5 commented that “The chart creation process is very user-friendly for me.
-There are many expressive templates available for me to use … Previously, I had to spend a lot of time on creating visualizations.”
+E1 mentioned that “This system can complete highly professional tasks without requiring extensive skills or knowledge." E5 noted that “the chart can be translated and rotated in 3D, and it integrates seamlessly with the video.” They were also satisfied with the chart creation aspect.
+E4 noted that “the data mapping feature eliminated the need for selecting data and creating visualizations in 3D software." E5 commented that “The chart creation process is very user-friendly for me.
+There are many expressive templates available for me to use … Previously, I had to spend a lot of time on creating visualizations."
 7 LIMITATIONS AND FUTURE WORK
-Despite the encouraging evaluation results, several discussions and limitations were identified during the design and implementation process or mentioned by the experts during interviews.
+While the evaluation results indicate ChartBlenderis promising to help users embed and synchronize charts with video scenes, the system still has several limitations that were found during the implementation or mentioned by the participants during interviews.
 We hope to guide potential future research directions by pointing out these limitations.
@@
 Future work may explore fine-tuning Large Language Models (LLMs) to support automatic topic analysis, enabling the system to generate context-aware visualizations with optimized design elements, enhancing the coherence and quality of data-driven videos.
-Layered Rendering.
-Although the current system supports depth-aware relationships between charts, Several experts (E1, E3) highlighted the potential of masking functionality that enables visualizations to be occluded by foreground objects in the video.
-Such a feature would allow charts to appear partially or fully behind scene elements, enhancing realism and spatial coherence.
 Tracking Editing.
@@
 This would significantly enhance the flexibility and expressiveness of the authoring workflow.
+Semantic Motion Analysis.
+While our system employs robust pixel‑level tracking to ensure precise geometric synchronization, it lacks the ability to handle occlusions effectively (E1,E2,E3).
+This limitation stems from its reliance on low‑level visual features without integrating higher‑level semantic reasoning.
+We regard pixel‑level tracking as a fundamental requirement for stable embedding, as it generalizes well across diverse video types without being constrained by predefined object categories.
+Introducing semantic recognition could effectively address motion‑occlusion challenges, enabling the system to maintain tracking continuity even when visual features are temporarily obscured.
 8 Conclusion
@@
 In addition, the system incorporates an interactive workflow to support easy and efficient authoring of videos with visualizations embedded.
-Quantitative evaluations demonstrate the effectiveness of our tracking algorithms in recovering spatially consistent motion trajectories.
+Quantitative evaluations demonstrate the effectiveness of our algorithm pipeline in recovering spatially consistent motion trajectories.
 Additionally, interviews with domain experts confirm the system’s usability, expressiveness, and potential for professional media production.
@@
 We would like to thank all the reviewers for their valuable feedback.
+[Uncaptioned image] Yi He received her B.Eng degree from School of Digital Media & Design Arts, Beijing University of Posts and Telecommunications.
+She is currently working toward her Ph.D.
+degree as part of the Intelligent Big Data Visualization (iDVx) Lab, Tongji University.
+Her research interests include data visualization and artificial intelligence.
+[Uncaptioned image] Yuqi Liu received her B.Eng degree from the School of Computer Science and Technology, Shanghai University of Computer Engineering and Science.
+She is currently working toward her Master’s degree as part of the Intelligent Big Data Visualization (iDVx) Lab at Tongji University.
+Her research interests include data visualization and artificial intelligence.
+[Uncaptioned image] Chenpu Li is pursuing his B.Eng dual degree in Visual Communication Design and Artificial Intelligence at the College of Design and Innovation, Tongji University.
+His cur-rent academic focus spans Human-Computer Interaction (CHI) and AI-Art integration.
+[Uncaptioned image] Ruoyan Chen received her B.Eng degree in Industrial Design from University of Science and Technology Beijing.
+She is currently pursuing her M.S.
+degree as part of the Intelligent Big Data Visualization (iDV*) Lab, Tongji University.
+Her research interests include data visualization and agent design.
+[Uncaptioned image] Chuer Chen received her MSc degree from the Department of Electrical and Computer Engineering, National University of Singapore in 2021.
+She is currently working toward her Ph.D.
+degree as part of the Intelligent Big Data Visualization (iDVx) Lab, Tongji University.
+Her research interests include information visualization and intelligent design.
+[Uncaptioned image] Shengqi Dang received his B.Sc.
+in Mathematics from Tongji University and is currently pursuing a Ph.D.
+degree at the Intelligent Big Data Visualization (iDVx) Lab, Tongji University.
+His research interests include of artificial intelligence and computer graphics.
+[Uncaptioned image] Nan Cao received his Ph.D.
+degree in Computer Science and Engineering from the Hong Kong University of Science and Technology (HKUST), Hong Kong, China in 2012.
+He is currently a professor at Tongji University and the Assistant Dean of the Tongji College of Design and Innovation.
+He also directs the Tongji Intelligent Big Data Visualization Lab (iDVx Lab) and conducts interdisciplinary research across multiple fields, including data visualization, human computer interaction, machine learning, and data mining.
+He was a research staff member at the IBM T.J.
+Watson Research Center, New York, NY, USA before joining the Tongji faculty in 2016.